### v1

# Quantum Electrodynamics of graphene Landau levels in a deep-sub-wavelength hyperbolic phonon polariton cavity

Abstract
The confinement of electromagnetic radiation within extremely small volumes offers an effective means to significantly enhance light-matter interactions, to the extent that zero-point quantum vacuum fluctuations can influence and control the properties of materials. Here, we develop a theoretical framework for the quantum electrodynamics of graphene Landau levels embedded in a deep-sub-wavelength hyperbolic cavity, where light is confined into ultrasmall mode volumes. By studying the spectrum, we discuss the emergence of polaritons, and disentangle the contributions of resonant quantum vacuum effects from those of purely electrostatic interactions. Finally, we study the hybridization between magnetoplasmons and the cavity’s electromagnetic modes.
I Introduction
Human beings naturally use optical light to perceive the surrounding world through sight. Leveraging technological advancements, other light frequencies can be used to probe the properties of matter. For example, X-rays reveal the atomic structure of crystals Compton , and Raman spectroscopy Raman measures vibrational modes of molecules. After the seminal work of Purcell Purcell1946 , it is now well established that light can be tailored by means of resonant structures - or cavities - to enhance its interaction with matter. Indeed, the emergence of cavity quantum electrodynamics (C-QED) Raimond01 ; Haroche13 , where mirrors confine the electromagnetic field, has enabled the measurement of quantum properties of a single atom coupled to a single photon.
Recently, a new paradigm has emerged: light can also be used to control matter cavityqm ; BlochReview , altering electronic structures and inducing novel states of matter. As a paradigmatic example, cold atoms in driven optical cavities exhibit altered many-body out-of-equilibrium properties, leading to emergent phases Ritsch2013 ; Baumann2010 . Remarkably, due to vacuum fluctuations, even an empty, non-driven cavity VacuumReview can change the properties of a material. Although only a few pioneering experiments have confirmed this concept Jarc2023 ; appugliese2022 ; enkner2024 , a growing body of theoretical predictions suggests that vacuum fluctuations could be harnessed to control the properties of matter Lu2024 ; schlawin2019 ; nguyen2023 ; wang2019 ; gao2020 ; Mazza2019 ; Sentef2018 leading to new cavity-induced mesoscopic phenomena, such as ferroelectricityLatini2021 and enhanced superconductivityKozin2024 . These methods generally necessitate advanced control over solid-state systems, such as two-dimensional (2D) electron gases subject to a quantizing magnetic field appugliese2022 ; enkner2024 .
Light-matter hybridization manifests as an avoided crossing between energy levels, known as the vacuum Rabi splitting $\Omega$Sanchez-Mondragon1983 ; agarwal1984 ; agarwal1985 ; Brune1996 ; Raimond01 , which in solid-state systems can be probed in transmission and reflection experimentsKhitrova2006 ; Todorov_PhysRevLett_2010 ; Todorov_PhysRevLett_2009 ; Scalari_Science_2012 ; Bitton . This phenomenon can be understood within both classical and quantum settingsDDB_arxiv_2023 but, in the context of C-QED, the quantum nature of the electric field $\hat{E}$ unveils a general correspondence between the vacuum Rabi splitting and quantum vacuum fluctuations of the electric field, $\Omega\sim\sqrt{\langle\hat{E}^{2}\rangle}$ dalibard1982 ; ciuti2005 ; DDB_arxiv_2023 . A promising strategy to enhance vacuum fluctuations (and consequently the coupling) is to confine the electromagnetic field to a small volume $V$chang2006 ; Kockum19 ; tame2013 , as the amplitude of fluctuations scales as $1/V$. While Fabry-Pérot cavities are often usedRaimond01 ; Haroche13 , the minimal achievable volume in these objects is limited by the so-called diffraction limit, $V\geq\lambda_{\rm d}^{3}$, where $\lambda_{\rm d}\equiv\lambda_{0}/2$ and $\lambda_{0}$ is the wavelength in free space. This limit imposes a significant obstacle to the compression and impact of quantum vacuum fluctuations.
Conversely, sub-wavelength cavities, made from an optically active electromagnetic environment, differ from optical cavities as they can confine light in a volume $V\ll\lambda_{\rm d}^{3}$, well below the diffraction limit chang2006 ; koppens2011 ; tame2013 ; bozhevolnyi2006 ; oulton2009 ; baumberg2019 ; tame2013 . Paradigmatic examples of sub-wavelength cavities are plasmonic systems and metamaterials Maier2007 . Plasmonic nanocavities exploit the excitation of surface plasmons at metal-dielectric interfaces to confine electromagnetic fields to volumes significantly smaller than the diffraction limit Schuller2010 ; Ciraci2012 ; Chikkaraddy2016 .
While these systems can achieve extremely high compression factorsAlcaraz2018 ; Barra-Burillo2021 , they suffer from a low quality factor ($Q\approx 10$) stemming from the highly dissipative nature of metallic systems. Among these, split-ring resonators operating in the terahertz regimeMaissen2014 have been extensively used in solid-state cavities experiments, due to their integrability with high-quality two-dimensional electron gases. More generally, polaritonic systems serve as promising platforms for solid-state C-QED, as they can host deep sub-wavelength optical modes and offer great tunability dai2014 . While this approach is experimentally promising, most current theories are developed for Fabry-Pérot cavities, where excitations are described by a transverse vector potential $\hat{\bm{A}}$. These theories are applied to sub-wavelength cavities by using a phenomenological substitution to enhance the Rabi frequency with a compression factor $A=V/\lambda_{\rm d}^{3}$. The compression factor quantifies how much a sub-wavelength cavity surpasses the diffraction limit Keller2017 ; arwas2023 , reaching values of $A\approx 10^{-10}$. Given a Rabi frequency $\Omega$ derived from optical cavity theory, which scales as $\sim 1/\sqrt{V}$, the same quantity for a sub-wavelength cavity is obtained by substituting $\Omega\to\Omega/\sqrt{A}$, thus enhancing light-matter interactionsgao2020 ; schlawin2019 ; wang2019 ; nguyen2023 ; zezyulin2022 ; lenk2022 ; arwas2023 ; pavosevic2022 by a factor up to $10^{5}$. This phenomenological procedure underlies striking predictions in sub-wavelength cavities, such as cavity-mediated superconductivity schlawin2019 , breakdown of topological integer quantum Hall effect arwas2023 , and cavity control of topological properties wang2019 ; nguyen2023 . However, it has been shown andolina2024 that sub-wavelength cavities, characterized by a strong compression $A\ll 1$, cannot be described by a transverse vector potential $\hat{\bm{A}}$ but, rather, by a scalar potential $\hat{\phi}$, reflecting their electrostatic characterandolina2024 ; pantazopoulos2024 . This underscores the need for a comprehensive theory of sub-wavelength cavity QED based on the coupling with a scalar potential $\hat{\phi}$. Consequently, it is important to revisit the conclusions of Refs. gao2020, ; schlawin2019, ; wang2019, ; nguyen2023, ; zezyulin2022, ; lenk2022, ; arwas2023, ; pavosevic2022, in light of this new framework. Developing this theory is the main goal of the present article.
The excitations of a sub-wavelength cavity are not transverse photons but longitudinal hybrid light-matter excitations such as plasmonsGPlasmonic and polaritonsPCavities . Sub-wavelength cavities have recently been shown riolo2024 to alter the quasiparticle properties of Fermi liquids realized in weakly correlated 2D electron systems.
Among sub-wavelength platforms, hyperbolic phonon-polariton (HPP) cavities dai2014 ; caldwell2014 ; HerzigSheinfux ; li2015 ; Autore based on hexagonal boron nitride (hBN) offer an unparalleled balance of high compression factor ($A\approx 10^{-6}$) and impressive quality factors ($Q\approx 400$) HerzigSheinfux . Due to the strong anisotropy in its optical phonon spectrumcaldwell2014 , hBN exhibits an upper ($\sim 168-200$ meV) and a lower ($\sim 94-103$ meV) Reststrhalen band, where the in-plane and out-of-plane permittivity components are opposite in sign, leading to a hyperbolic polaritonic dispersion. The resulting ray-like nature of hBN mid-infrared phonon polaritons allows for extreme wavelength compression of light, enabling electromagnetic waves to be confined to deeply sub-wavelength scalesli2015 . On the other hand, the bulk nature of hBN phonon polaritons protects them from Ohmic loss mechanisms typical of metal-dielectric interfaces, enhancing the lifetime and making HPP cavities highly effective for mid-IR photonic applicationscaldwell2014 .
In this paper, we present a fully non-local quantum theory to describe the cavity quantum electrodynamics of 2D electron systems coupled to sub-wavelength cavities, aiming to derive the vacuum Rabi frequencies from a fully microscopic theoretical framework without relying on any phenomenological assumptions. Specifically, we investigate electrons in graphene subject to a strong perpendicular magnetic field ${\bm{B}}$ and coupled to the sub-wavelength modes of HPP cavity. This ${\bm{B}}$ field induces circular electron motion (cyclotron motion), leading to the emergence of quantized energy levels known as Landau levels (LLs)Giuliani_and_Vignale ; PrangeGirvin1987 . The excitations of HPP cavities made from hBN are typically around $0.1~{}{\rm eV}$ HerzigSheinfux , hence LL transitions can be tuned to this energy range by applying a magnetic field of a few Tesla. Therefore, if the coupling is higher than the losses, LLs transitions hybridize with resonant cavity modes, forming Landau-phonon-polaritons (LPhPs) hagenmuller2010 ; hagenmuller2012 ; Chirolli2012 ; Pellegrino2014 ; ashida2023 .
To describe LPhPs, we develop a comprehensive microscopic model of quantum vacuum fluctuations in deep sub-wavelength HPP cavities. With our microscopic theory, we can predict the optimal cavity size to maximize light-matter coupling, measured by the vacuum Rabi frequencies, without resorting to any phenomenological compression factor. Importantly, the deep sub-wavelength regime allows cavity and matter excitations to hybridize at very high wavevectors, necessitating a fully non-local theoretical framework. Indeed, the maximal vacuum Rabi frequencies are obtained for finite momenta. Scattering-type near-field optical microscopy (s-SNOM) basov2016 , capable of probing collective modes at such high wavevectors, provides a promising platform to measure our predictions. The fully microscopic theory developed here can have profound implications for the control of materials in sub-wavelength cavities. As a major example, it will be interesting in the future to use our theory to study quantum transport in these 2D electron systems embedded in HPP cavities. Several experiments showed that sub-wavelength split-ring cavities affect the integer appugliese2022 ; enkner2024 and fractional enkner2024b quantum Hall, generating significant theoretical interest ciuti2021 ; arwas2023 ; bacciconi2023 ; rokaj2023 ; rokaj2022 . In principle, our framework can be applied to study how HPP cavities modify magnetotransport paravicini2019 and topology in general arwas2023 ; nguyen2023 ; wang2019 ; Zeno ; Dmytruk2022 ; Dmytruk2024 ; Dag2024 . In our case, as the collective magnetic excitations are strongly coupled to the cavity, a fully many-body treatment of quantum transport would be needed, rather than a single-particle approach. Since in sub-wavelength cavities the light-matter coupling occurs via the density operator, rather than via the current, these cavities circumvent the so-called “no-go theorems”andolina_prb_2019 ; andolina_epjp_2022 ; rzazewski_prl_1975 ; nataf_naturecommun_2010 and thus can, in principle, realize a superradiant phase transition. However, since the excitations of the cavity are encoded in the scalar potential $\hat{\phi}$, the nature of the phase transition would be ferroelectricdebernardis2018 ; rouse2023 , and would not lead to photon condensation.
This paper is organized as follows. In Section II, we summarize the main concepts and key results of this work, which are presented in greater detail in the subsequent sections. In Section III, we develop a quantum theory of HPP cavities. Section IV provides an overview of the physics of LLs in graphene. In Section V, we delve into the details of light-matter interactions and analyze the resulting polaritonic spectrum. Additional technical details regarding the electrostatic properties of hexagonal boron nitride are provided in Appendix A, while the density-density response function of Landau levels in graphene is reviewed in Appendix B.
II Summary of main results
In this work, we present our primary findings structured as follows: First, in Sect. II.1, we solve the retarded Poisson equation within a sub-wavelength cavity to determine the scalar Green’s function. This solution enables us to quantize the scalar potential $\hat{\phi}$, expressing it in terms of the quantum modes of the cavity. This provides a theoretical framework for describing electromagnetic and quantum vacuum fluctuations within a subwavelength cavity. Thus, in Sect. II.2 we review the physics of LLs emerging in graphene subjected to a strong perpendicular magnetic field. Crucially, we highlight that, in sub-wavelength cavities, electrons couple through the density rather than the current as in traditional Fabry-Pérot cavities. Finally, in Sect. II.3 we develop the C-QED of a 2D material embedded in a sub-wavelength cavity. By formulating and diagonalizing the Hopfield Hamiltonian, we study polaritonic states that arise from the hybridization of LL transitions with cavity modes. This also allows us to directly calculate, without invoking any phenomenological procedure, the vacuum Rabi frequencies of the system, showing the system achieves ultra-strong and super-strong coupling,
While our theory is specifically applied to graphene embedded in an hBN cavity, the same theory can be readily applied to other 2D materials within sub-wavelength cavities.
II.1 Theory of the Green’s function of HPP cavities
[…]
Figure 1: (Color online) A schematic representation of the system under study. The HPP cavity is formed by two metallic mirrors (depicted in gold) located at $|z|>L_{z}/2$. The region between the mirrors is filled with hBN, shown in light blue. A graphene sheet is positioned at $z=0$ and subjected to an external perpendicular magnetic field.
Generally, electromagnetic cavities are fully characterized by the solutions of Maxwell’s equationsJackson . In sub-wavelength cavities the most important contribution to the description of the electromagnetic properties is provided by the solution of the retarded Poisson equation in the presence of an external chargeandolina2024 ; riolo2024 :
$-\partial_{i}[\epsilon_{ij}(\bm{r},\omega)\partial_{j}\phi(\bm{r},\omega)]=4\pi\rho_{\rm ext}(\bm{r},\omega)~{},$ (1)
where $\bm{r}$ is the position, $\partial_{i}$ is a shorthand for the spatial derivative $\partial/\partial r_{i}$, $\omega$ is the frequency, $\epsilon_{ij}(\bm{r},\omega)$ is the dielectric tensor of the cavity embedding, $\phi(\bm{r},\omega)$ is the potential, and $\rho_{\rm ext}(\bm{r},\omega)$ is the free charge. In the case of sub-wavelength cavities, the speed of light $c$ can be considered infinite, since $\lambda\ll\lambda_{0}$. Hence, we restrict our study to the electrostatic component. Note, however, that we fully account for the retardation associated with the response of the metallo-dielectric environment creating the cavity, which is encoded in the permittivity tensor $\epsilon_{ij}(\bm{r},\omega)$. Consequently, while the electromagnetic behavior is governed by the electrostatic Poisson equation (Eq. (1)), the dynamic behavior of the matter degrees of freedom is thoroughly taken into account.
The cavity geometry that we consider is shown in Fig. 1. We further assume translational invariance along the $\hat{\bm{x}}$ and $\hat{\bm{y}}$ directions that denote 2D positions in this plane with the vector $\bm{r}_{\parallel}=x\hat{\bm{x}}+y\hat{\bm{y}}$. In the direction perpendicular to this plane, i.e. the $\hat{\bm{z}}$ direction and for $|z|\leq L_{z}/2$, we assume that space is filled by a hBN layer of thickness $L_{z}$ dai2014 ; caldwell2014 ; HerzigSheinfux ; li2015 ; Autore , characterized by a uniaxial permittivity tensortomadin2015 ; riolo2024 $\bm{\epsilon}(\bm{r},\omega)=\text{diag}(\epsilon_{x}(\bm{r},\omega),\epsilon_{x}(\bm{r},\omega),\epsilon_{z}(\bm{r},\omega))$. This dielectric tensor is initially assumed to be real, with its imaginary part introduced later through the substitution $\omega\to\omega+\mathrm{i}\eta$. While this approach is not generally required, it is particularly advantageous in this context, as it allows us to quantize the field first and incorporate losses subsequently.
Further details on this tensor are reported in Appendix A. For $|z|>L_{z}/2$, we assume the presence of two perfect conductors, acting as mirrors. Consequently, the boundary conditions for the potential inside hBN ( $|z|\leq L_{z}/2$) are given by Jackson : $\phi(\bm{r}_{\parallel},z=\pm L_{z}/2)=0$.
[…]
Figure 2: (Color online) Different spectra are plotted as functions of $q_{\parallel}$ (expressed in ${\rm nm}^{-1}$) and $\hbar\omega$ (expressed in $\rm{eV}$). Panel (a) shows the imaginary part of the bare Green’s function of the HPP cavity, i.e. $-e^{2}{\rm Im}~{}g^{(0)}(q_{\parallel},0,0,\omega)$, for $L_{z}=50~{}{\rm nm}$ and taking a cutoff on the modes, $n_{z}^{\rm max}=50$. Panel (b) displays the imaginary part of the bare density-density response function for LLs in graphene, $-{\rm Im}~{}\chi^{(0)}_{\rm nn}({q}_{\parallel},\omega)$, taking $B=5~{}{\rm T}$ and a cutoff for the LLs $N_{\rm c}=30$. In both panels we choose $\hbar\eta=5\times 10^{-4}{\rm meV}$.
Instead of directly solving Eq. (1) for an arbitrary external charge distribution, it is convenient to solve for the associated scalar bare Green’s functionandolina2024 $g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)$ corresponding to the solution of Eq. (1) in the presence of a test charge $\rho^{({\rm test},\bm{r}^{\prime})}_{\rm ext}(\bm{r},\omega)=\delta(\bm{r}-\bm{r}^{\prime})$ at a given position $\bm{r}^{\prime}$. Note that the bare Green’s function $g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)$ describes how an empty HPP cavity responds to a test charge, while the effect of an embedded material will be treated later in Section II.3. Once the Green’s function has been obtained, the potential $\phi(\bm{r},\omega)$ can be obtained directly as a convolution between $g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)$ and the external density $\rho_{\rm ext}(\bm{r}^{\prime},\omega)$. The scalar Green’s function of an HPP cavity can be obtained from the following Fourier transform with respect to the 2D coordinate ${\bm{r}}_{\|}$ and expansion in normal modes:
$\begin{split}&g^{(0)}(\bm{q}_{\parallel},z,z^{\prime},\omega)=\frac{8\pi}{L_{z}}\sum_{n_{z}=0}^{\infty}\frac{\cos\left(q_{z,n_{z}}z\right)\cos\left(q_{z,n_{z}}z^{\prime}\right)}{{\epsilon}_{z}(\omega)q_{z,n_{z}}^{2}+{\epsilon}_{x}(\omega)q_{\parallel}^{2}}~{},\end{split}$ (2)
where $\bm{q}_{\parallel}$ is the wavevector in the $\hat{\bm{x}}-\hat{\bm{y}}$ plane, ${q}_{\parallel}$ is its modulus, and $q_{z,n_{z}}\equiv(2n_{z}+1)\pi/L_{z}$. Eq. (LABEL:eqn:green_function_hBN_expansion_Main) represents a key theoretical result, as it shows that the scalar Green’s function can be expanded into distinct modes, which can then be quantized systematically. Notice that the Green’s function depends only on the modulus ${q}_{\parallel}$ as the system is invariant under rotations in the $\hat{\bm{x}}-\hat{\bm{y}}$ plane. Fig. 2(a) displays the spectrum $-{\rm Im}~{}g^{(0)}({q}_{\parallel},0,0,\omega)$ of the HPP cavity, which quantifies the absorption of the cavity at the position $z=z^{\prime}=0$, where we will place the 2D material. The scalar Green’s function $g^{(0)}({q}_{\parallel},z,z^{\prime},\omega)$ exhibits poles when the condition ${\epsilon}_{z}(\omega)q_{z,n_{z}}^{2}+{\epsilon}_{x}(\omega)q_{\parallel}^{2}=0$ is satisfied. This occurs only in the lower ($\omega_{z}^{\rm T}<\omega<\omega_{z}^{\rm L}$) and upper ($\omega_{x}^{\rm T}<\omega<\omega_{x}^{\rm L}$) Reststrahlen bands. The longitudinal and transverse phonon frequencies, $\omega_{i}^{\rm L}$ and $\omega_{i}^{\rm T}$, are marked by black dotted lines in Fig. 2(a). We denote the poles of the bare scalar Green’s function with the symbol $\omega_{{q}_{\parallel},n_{z},s}$, where $s=+1$ ($s=-1$) refers to the upper (lower) Reststrahlen band. There are important differences and similarities between the scalar Green’s function in Eq. (LABEL:eqn:green_function_hBN_expansion_Main) for HPP cavities and the full dyadic Green’s function for Fabry-Pérot cavities andolina2024 ; Charlie_Thesis . Both functions share key features, such as poles corresponding to mode frequencies and a prefactor of $1/L_{z}$, which encodes the compression mechanism. However, in both cases, the dependence on $L_{z}$ is non-trivial, as the mode frequencies $\omega_{{q}_{\parallel},n_{z},s}$ themselves depend on the cavity length $L_{z}$. In Fabry-P’erot cavities, the poles, which are associated with the energies of cavity photons, appear only in the vectorial part of the dyadic Green’s function andolina2024 , while the scalar part only represents the Coulomb potential modified by the mirrors’image charges. Conversely, in HPP cavities, poles corresponding to the phonon-polariton energies are present in both the scalar and vectorial components of the dyadic Green’s function. Nevertheless, due to the sub-wavelength nature of HPP cavity modes ($\lambda_{\rm p}=2\pi/q_{\parallel}\ll\lambda_{0}$) the scalar Green’s function $g^{(0)}({q}_{\parallel},z,z^{\prime},\omega)$ is the dominant contribution to the system’s electromagnetic properties andolina2024 .
Note that, in the static limit ($\omega=0$), the scalar Green’s function reduces to the Coulomb potential, as screened by the metallic mirrors at $z=\pm L_{z}/2$ and the static polarization charges inside hBN. Taking also the limit $L_{z}\to\infty$, $g^{(0)}({q}_{\parallel},0,0,0)$ reduces to the 2D Coulomb potential, as renormalized by the static dielectric properties of hBN, i.e. $\lim_{L_{z}\to\infty}g^{(0)}({q}_{\parallel},0,0,0)=2\pi/(q_{\parallel}\sqrt{\epsilon_{x}(0)\epsilon_{z}(0)})$. We conclude that Eq. (LABEL:eqn:green_function_hBN_expansion_Main) accurately describes both the static screening ($\omega=0$) induced by hBN polarization charges and metallic mirrors, and the resonant structure at finite $\omega$, occuring at the location of the poles $\omega_{{q}_{\parallel},n_{z},s}$.
In free space, the Poisson equation in Eq. (1) lacks a time-dependent component, making $\phi(\bm{r})$ a completely classical field. However, in HPP cavities, the non-trivial frequency dependence of the dielectric permittivity tensor causes the emergence of resonant modesGubbin_JChemPhys_2022 , which can be described as quantum mechanical modes. Consequently, the poles of Eq. (LABEL:eqn:green_function_hBN_expansion_Main) correspond to the quantum mechanical modes of the cavity Hamiltonian, ^$\hat{\mathcal{H}}_{\rm cav}=\sum_{\bm{q}_{\parallel},n_{z},s}\hbar\omega_{{q}_{\parallel},n_{z},s}\hat{a}^{\dagger}_{\bm{q}_{\parallel},n_{z},s}\hat{a}_{\bm{q}_{\parallel},n_{z},s}$, where the operator $\hat{a}^{\dagger}_{\bm{q}_{\parallel},n_{z},s}$ ($\hat{a}_{\bm{q}_{\parallel},n_{z},s}$) creates (destroys) a cavity excitation. Since the cavity is a deeply sub-wavelength one, the electromagnetic field is well described by a scalar potential
$\hat{\phi}(\bm{r}_{\parallel},z)$ $=\sum_{\bm{q}_{\parallel},n_{z},s}{\phi_{{q}_{\parallel},n_{z},s}(z)}\left(\frac{e^{-i\bm{q}_{\parallel}\cdot\bm{r}_{\parallel}}}{{\sqrt{S}}}\hat{a}^{\dagger}_{\bm{q}_{\parallel},n_{z},s}+{\rm h.c.}\right)~{},$ (3)
where $S$ is the surface area of the cavity, and $\phi_{{q}_{\parallel},n_{z},s}(z)\equiv\cos(q_{z,n_{z}}z)\sqrt{{2}/({N_{{q}_{\parallel},n_{z},s}L_{z}})}$, with $N_{{q}_{\parallel},n_{z},s}$ being a normalization factor. The procedure to determine this normalization factor will be detailed momentarily. The physical electric field can be obtained by taking the gradient of the potential in Eq. (3). Further details are reported in Subsection III.1. To determine this normalization factor, we employ the following strategy. First, we calculate the quantum scalar Green’s function $g_{\rm q}^{(0)}({q}_{\parallel},z,z^{\prime},\omega)$ using the scalar potential $\hat{\phi}(\bm{r}_{\parallel},z)$ in Eq. (3) and the standard Kubo formulaGiuliani_and_Vignale as detailed in Subsection III.1. As we will show momentarily, the scalar potential $\hat{\phi}(\bm{r}_{\parallel},z)$ expanded in terms of the HPP modes $\hat{a}_{\bm{q}_{\parallel},n_{z},s},\hat{a}^{\dagger}_{\bm{q}_{\parallel},n_{z},s}$ in Eq. (3) does not describe the full electrostatic contribution. Instead, there is a frequency-independent contribution that is not captured by these modes. As only resonant processes contribute to dissipation, all dissipative processes are accounted by $g_{\rm q}^{(0)}({q}_{\parallel},z,z^{\prime},\omega)$, the imaginary part of the quantum Green’s function matches that of the Green’s function derived from the Poisson equation as given in Eq. (LABEL:eqn:green_function_hBN_expansion_Main) i.e. ${\rm Im}~{}g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)={\rm Im}~{}g^{(0)}(q_{\parallel},z,z^{\prime},\omega)$. By imposing this condition we determine the normalization factor $N_{{q}_{\parallel},n_{z},s}$ by imposing that ${\rm Im}~{}g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)={\rm Im}~{}g^{(0)}(q_{\parallel},z,z^{\prime},\omega)$. Still, while the imaginary parts coincide, the real parts of $g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)$ and $g^{(0)}(q_{\parallel},z,z^{\prime},\omega)$ differ.
Specifically, $g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)$ vanishes in the high-frequency limit, whereas $g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})\equiv\lim_{\omega\to\infty}g^{(0)}(q_{\parallel},z,z^{\prime},\omega)$ remains finite. As a result, the quantum Green’s function $g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)$, associated with the modes $\hat{a}_{\bm{q}_{\parallel},n_{z},s}$ and $\hat{a}^{\dagger}_{\bm{q}_{\parallel},n_{z},s}$, satisfies the Kramers-Kronig relations. In contrast, the full scalar Green’s function $g^{(0)}(q_{\parallel},z,z^{\prime},\omega)$ is not analytic on the complex plane, as $g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})$ is nonzero. Consequently, $g^{(0)}(q_{\parallel},z,z^{\prime},\omega)$ fulfills the Kramers-Kronig relations only when $g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})$ is subtracted. Indeed, the following decomposition holds: $g^{(0)}(q_{\parallel},z,z^{\prime},\omega)=g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)+g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})$. The term $g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)$ effectively captures the resonant processes tied to the excitation of the system’s quantum mechanical modes $\hat{a}_{\bm{q}_{\parallel},n_{z},s},\hat{a}^{\dagger}_{\bm{q}_{\parallel},n_{z},s}$, whereas $g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})$ accounts for a frequency-independent renormalization of the Coulomb interaction. The details of this section are provided in Sec. III.
[…]
Figure 3: (Color online) The RPA density-density response of the coupled system (HPP cavity plus graphene), $-{\rm Im}~{}{\chi}_{\rm nn}({q}_{\parallel},\omega)$. Panel (a) shows $-{\rm Im}~{}{\chi}_{\rm nn}({q}_{\parallel},\omega)$ as a function of the wavevector $q_{\parallel}$ for a given magnetic field, $B=5~{}{\rm T}$ (corresponding to $\ell_{B}\approx 11.5~{}{\rm nm}$). Panel (b) shows the same quantity as a function of $\hbar v_{\rm F}/\ell_{B}$, for fixed momentum $\bar{q}_{\parallel}=0.1~{}{\rm nm}^{-1}$. Other parameters are $L_{z}=50~{}{\rm nm}$, $N_{\rm c}=30$, $n_{z}^{\rm max}=30$, and $\hbar\eta=5\times 10^{-4}~{}{\rm meV}$.
II.2 Theory of Landau levels in Graphene and their coupling to a subwavelength cavity
We now consider single-layer graphene, which will be later positioned at the center of the HPP cavity (i.e. at $z=0$), subject to a strong, uniform, and static perpendicular magnetic field $B$. In this regime, electronic excitations are inter-Landau-level transitions Goerbig11 ; ando1982 ; roldan2009 ; roldan2010 . In the Landau gauge, the energy spectrum of electrons roaming in this crystal are massless Dirac fermion Landau levels (LLs) given by $\varepsilon_{\lambda,n}=\lambda(\hbar v_{\rm F}/\ell_{B})\sqrt{2n}$, where $\ell_{B}=\sqrt{\hbar c/(eB)}$ is the magnetic length, $e$ is the electron charge, $v_{\rm F}$ is the Fermi velocity in graphene, $k_{x}$ is the electronic wavevector along the $x$-axis, which accounts for the degeneracy of the spectrum, $n$ is a non-negative integer representing the LL index, and $\lambda=\pm$ denotes conduction/valence band states. We denote by the symbol $N_{\rm F}$ the index of the last occupied LL. Here, we concentrate on the charge neutrality point of graphene, choosing $N_{\rm F}=0$. In Fig. 2(b), we plot the non-interacting spectrum of LLs, $-{\rm Im}\chi^{(0)}_{\rm nn}({q}_{\parallel},\omega)$, where $\chi^{(0)}_{\rm nn}({q}_{\parallel},\omega)$ is the bare density-density response function calculated using the Kubo formula roldan2009 ; Giuliani_and_Vignale . This quantity is fully non-local, valid for all wavevectors ${q}_{\parallel}$, which is particularly relevant as the deep sub-wavelength nature of hBN enables exploration of high wave-vector regions. Specifically, accounting for spatial non-local effects is necessary in the limit of rapidly varying matter fields ($q\sim 0.1~{}{\rm nm}^{-1}$), which has been investigated in SNOM experimentsdai2014 , and also in the presence of a magnetic fieldwehmeier2024 .
The LLs spectrum shows electron-hole pair excitations associated with transitions $\lambda,k_{x},n\to\lambda^{\prime},k_{x}^{\prime},n^{\prime}$ and corresponding transition energies $\varepsilon_{\lambda^{\prime},n^{\prime}}-\varepsilon_{\lambda,n}$. Due to the $k_{x}$ degeneracy, bare LL transitions do not depend on the momentum transfer ${q}_{\parallel}=k_{x}^{\prime}-k_{x}$ and thus appear as flat lines in the ${q}_{\parallel}-\omega$ plane. In Fig. 2(b), we plot transitions between valence-band states $\lambda=-,n$ and conduction-band states $\lambda^{\prime}=+,n^{\prime}$. Dipole-allowed transitions $n^{\prime}=n\pm 1$ are drawn as solid white lines, while other transitions are shown as white dashed lines. In the numerics, we choose a strong magnetic field in the order of $B\approx 5~{}{\rm T}$ (corresponding to $\ell_{B}\approx 11.5~{}{\rm nm}$). Due to the deep sub-wavelength nature of our cavities, we focus on high wavevectors $q_{\parallel}\approx 0.1~{}{\rm nm}^{-1}$, where all transitions are relevant, as $q_{\parallel}\ell_{B}\sim 1$. The discrete nature of LLs leads to a series of transition frequencies that can couple with the electromagnetic modes of the HPP cavity.
We now move one to discuss the coupling between matter degrees of freedom and cavity modes. In the Coulomb gauge, the LLs and the cavity modes interact through the electrostatic interaction Hamiltonian andolina2024 ; Giuliani_and_Vignale :
[…]
where $\hat{\phi}(\bm{r}_{\parallel},z=0)$ is the scalar potential of the cavity reported in Eq. (3) and evaluated at the vertical location of graphene and $\hat{n}(\bm{r}_{\parallel})$ is the electron density operator in graphene. This Hamiltonian describes the coupling of the electronic density in graphene with the electric field of the cavity, leading to the formation of hybrid light-matter states known as polaritons hopfield1958 ; Kavokin11 . However, the interaction Hamiltonian in Eq. (4) only accounts for the dynamical part of the field, corresponding to the quantum Green’s function $g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)$. Therefore, it is necessary to include the effect of the statically-screened potential, $g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})$. This can be incorporated into the Hamiltonian in the form of a standard two-body screened Coulomb interaction, ^$\hat{\mathcal{H}}_{\rm C}=({1}/{2S})\sum_{\bm{q}_{\parallel}}\hat{n}_{-\bm{q}_{\parallel}}e^{2}g_{\infty}^{(0)}(\bm{q}_{\parallel},0,0)\hat{n}_{\bm{q}_{\parallel}}$, where $\hat{n}_{\bm{q}_{\parallel}}$ is the momentum-resolved density operator. In the previous summation, the ${\bm{q}}_{\parallel}={\bm{0}}$ contribution is canceled by the inclusion of a positive background (Jellium) to ensure charge neutrality Giuliani_and_Vignale . Since our analysis focuses on finite $q_{\parallel}$, this procedure does not affect the results.
In sub-wavelength cavities, the interaction Hamiltonian involves a fundamentally different coupling compared to the traditional minimal coupling $\hat{\bm{p}}\to\hat{\bm{p}}+(e/c)\hat{\bm{A}}$ interaction, where $\hat{\bm{A}}$ is the transverse vector potential and $\hat{\bm{p}}$ is the momentum operator. The minimal coupling results into a diamagnetic term in the continuum-model Hamiltonian andolina_prb_2019 ; Ruggenthaler2023 ; Schaefer2020 or introduces a non-linear Peierl’s phase in truncated lattice models DeBernardis18 ; savasta2021gauge ; distefano2019 ; dmytruk2021 . In contrast, Eq. (4) represents a purely linear coupling, which is a distinctive feature of interactions within sub-wavelength cavities. The details of this section are provided in Sec. IV.
[…]
Figure 4: (Color online) The RPA proper density-density response of the coupled system (HPP cavity plus graphene), $-{\rm Im}~{}{\chi}_{\rm nn}(\bar{q}_{\parallel},\omega)$ as a function of $\hbar v_{\rm F}/\ell_{B}$ for a doped system ($N_{\rm F}=5$), for a fixed wavevector ($\bar{q}_{\parallel}=0.1{\rm nm}^{-1}$). Panel (a) illustrates this quantity $-{\rm Im}~{}{\chi}_{\rm nn}(\bar{q}_{\parallel},\omega)$ in the case where quantum mechanical modes are not included, $g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)=0$. Panel (b) shows the same quantity in the presence of these modes. Other parameters are $L_{z}=50{\rm nm}$, $N_{\rm c}=30$, $n_{z}^{\rm max}=50$, $\hbar\eta=5\times 10^{-4}{\rm meV}$. Panel (c) is the same of panel (b), with $L_{z}=10^{3}{\rm nm}$ and $n_{z}^{\rm max}=100$
II.3 The polaritonic spectrum and the vacuum Rabi frequency
Polaritonic excitations can be observed in the momentum-dependent absorption spectrum, $S_{\delta}({q}_{\parallel},\omega)$, where $\delta={\rm el}$ ($\delta={\rm cav}$) denotes the electronic (cavity) dynamic structure factor. The polaritonic spectrum is derived from the collective excitations of the coupled system. Within the Random Phase Approximation (RPA)Giuliani_and_Vignale , polaritons can be found by solving the equation for the poles of the dielectric functionriolo2024 $\epsilon_{\rm RPA}({q}_{\parallel},\omega)=0$. The dielectric function within this approximation is given byriolo2024 :
$\begin{split}\epsilon_{\rm RPA}({q}_{\parallel},\omega)\equiv 1-e^{2}g^{(0)}({q}_{\parallel},0,0,\omega)\chi^{(0)}_{\rm nn}({q}_{\parallel},\omega)~{}.\end{split}$ (5)
The zeros of $\epsilon_{\rm RPA}({q}_{\parallel},\omega)$ correspond to the resonant frequencies of the coupled system, indicating the formation of polaritons. Once $\epsilon_{\rm RPA}({q}_{\parallel},\omega)$ is known, we can calculate the dressed electronic spectrum using $S_{\rm el}({q}_{\parallel},\omega)=-(\hbar/\pi){\rm Im}~{}{\chi}_{\rm nn}({q}_{\parallel},\omega)$, where ${\chi}_{\rm nn}({q}_{\parallel},\omega)={\chi}^{(0)}_{\rm nn}({q}_{\parallel},\omega)/\epsilon_{\rm RPA}({q}_{\parallel},\omega)$ and the dressed cavity spectrum $S_{\rm cav}({q}_{\parallel},\omega)=-(\hbar/\pi){\rm Im}~{}{g}({q}_{\parallel},0,0,\omega)$, where ${g}({q}_{\parallel},0,0,\omega)={g}^{(0)}({q}_{\parallel},0,0,\omega)/\epsilon_{\rm RPA}({q}_{\parallel},\omega)$.
Alternatively, one can diagonalize the total Hamiltonian ^$\hat{\cal{H}}_{\rm tot}=\hat{\cal{H}}_{\rm LLs}+\hat{\cal{H}}_{\rm C}+\hat{\cal{H}}_{\rm cav}+\hat{\cal{H}}_{\rm int}$. In the single excitation manifold, electron-hole excitations between LLs can be treated as bosonic quasiparticles ciuti2005 . In this limit, the total Hamiltonian reduces to a quadratic bosonic Hopfield Hamiltonianhopfield1958 , which can be diagonalized exactly. The polariton energies obtained in this procedure are identical to those that one finds from the zeros of dynamical dielectric function $\epsilon_{\rm RPA}({q}_{\parallel},\omega)$.
Fig. 3 shows the electronic spectrum at the neutrality point $N_{\rm F}=0$. Panel (a) presents this spectrum as a function of the wavevector $q_{\parallel}$, while panel (b) displays the spectrum as a function of $(\hbar v_{\rm F}/\ell_{B})$. The latter spectrum can be experimentally obtained by varying the external magnetic field $B$ zhang2005 . Inside the Reststrahlen bands (defined by $\omega^{\rm T}_{i}<\omega<\omega^{\rm L}_{i}$, with these frequencies indicated by black dashed lines), the LL transitions hybridize with HPP cavity modes. The vacuum Rabi frequencies of the system represent the interaction strength between a selected LL transition and a cavity mode. As detailed in Section V, one of these couplings can reach up to $10\%$ of the cavity frequency, placing our system in the ultra-strong coupling regimeKockum19 ; geiser2012 . Furthermore, a single LL transition hybridizes with multiple cavity modes, a regime known as super-strong coupling kuzmin2019 . This region of the spectrum is mainly controlled by the LL transitions encoded in ${\chi}^{(0)}_{\rm nn}({q}_{\parallel},\omega)$ and by the quantum Green’s function associated with the cavity modes ${g}^{(0)}_{\rm q}({q}_{\parallel},0,0,\omega)$, corresponding to the expansion in Eq. (3). Outside the Reststrahlen bands, LL transitions experience a $q_{\parallel}$-dependent positive frequency shift, known as the depolarization shiftGiuliani_and_Vignale ; shtrichman2001 , leading to finite dispersion as apparent in Fig. 3(a). Panel (b) shows that, even off-resonance, the LL transitions do not coincide with the non-interacting case (indicated by white solid and dashed lines). This shift originates from the frequency-independent screening contribution of the Green’s function $g_{\infty}^{(0)}({q}_{\parallel},0,0)$ rather than from the HPP poles, as it occurs outside the Reststrahlen bands, where $\omega\neq\omega_{{q}_{\parallel},n_{z},s}$ This effect is largely due to the confinement induced by the two metallic mirrors. Indeed, this effect is significantly suppressed in the limit $L_{z}\to\infty$ (see Fig. 8(c) in Sect. V and the related discussion), which corresponds to removing the metallic mirrors and filling the entire space with hBN.
Fig. 4 shows the electronic spectrum for a doped system with $N_{\rm F}=5$, corresponding to a filling factor $\nu=22$, as $\nu=4N_{\rm F}+2$. At finite charge density $N_{\rm F}\neq 0$ and strong magnetic fields, a plasmonic excitation appears in the system, usually referred to as magnetoplasmon. Magnetoplasmons Giuliani_and_Vignale ; roldan2009 ; muravev2016 arise in doped semiconductors under a magnetic field due to the coupling between plasma oscillations of free carriers and their cyclotron resonance. In Fig. 4(a) we show the spectrum without the contribution of the quantum mechanical modes (which is obtained by setting $g_{\rm q}^{(0)}({q}_{\parallel},0,0,\omega)=0$), revealing the expected magnetoplasmon dispersionroldan2009 . Fig. 4(b) is obtained by taking into account the full cavity contribution. Here, the magnetoplasmon strongly hybridizes with the cavity modes achieving a Rabi spitting of $\approx 25\%$ of the cavity frequency. This significant coupling arises from the well-known fact that most of the oscillator strength, which quantifies the ability of a system to absorb or emit electromagnetic radiation, is associated with the magnetoplasmon resonanceGiuliani_and_Vignale . Finally, we note that in the $L_{z}\to\infty$ limit the Rabi splitting shown in Fig. 4(c) remains of a similar order. Although the coupling between LLs decreases with reduced compression, the corresponding increase in the density of cavity modes compensates, resulting in a splittingDai2015 ; Kumar2015 comparable to the case in which $L_{z}$ is finite. This effect is peculiar of subwavelength cavities, as transverse contributions from the vector potential $\hat{\bm{A}}$ vanish in the limit $L_{z}\to\infty$ ashida2023 . Similar considerations hold for the cavity spectrum, $S_{\rm cav}({q}_{\parallel},\omega)$. The details of this section are provided in Sec. V.
III Details on “Theory of the Green’s function of HPP cavities”
In this Section, we outline the key steps and methods used to derive the scalar Green’s function for HPP cavities. A quantum framework is presented to relate the Green’s function to cavity field quantization, with explicit expressions connecting the scalar potential’s vacuum fluctuations to mode coefficients. Finally, the static and dynamic components of the Green’s function are distinguished, ensuring a comprehensive description of both resonant and off-resonant effects in the cavity’s electromagnetic response. These results form the theoretical foundation for the study of hPP cavities and their interaction with embedded materials.
We now calculate the scalar Green’s function $g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)$ for the HPP cavity described above. This Green’s function is determined by the Poisson equation:
$-\partial_{j}^{\prime}[\epsilon_{ji}(\bm{r}^{\prime},\omega)\partial_{i}^{\prime}g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)]=4\pi\delta(\bm{r}-\bm{r}^{\prime})~{},$ (6)
where $\epsilon_{ji}(\bm{r}^{\prime},\omega)$ is the permittivity tensor given in Eq. (92), and $\partial_{j}^{\prime}$ denotes the partial derivative with respect to the coordinates of $\bm{r}^{\prime}$.
The electric potential $\phi(\bm{r},\omega)$ in the presence of an external charge $\rho_{\rm ext}(\bm{r}^{\prime},\omega)$ is given by the convolution of the Green’s function with the charge density:
[…]
We now lay down a generic mode expansion technique to solve Eq. (6). The Poisson equation can be rewritten as
$\hat{O}_{\bm{r}^{\prime}}(\omega)g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)=4\pi\delta(\bm{r}-\bm{r}^{\prime})~{},$ (8)
where $\hat{O}_{\bm{r}^{\prime}}(\omega)f(\bm{r}^{\prime},\omega)\equiv-\partial_{j}^{\prime}[\epsilon_{ji}(\bm{r}^{\prime},\omega)\partial_{i}^{\prime}f(\bm{r}^{\prime},\omega)]$ and $f(\bm{r}^{\prime},\omega)$ is an arbitrary function.
We introduce an orthonormal basis set $f_{\mu}(\bm{r},\omega)$ such that:
$O_{\bm{r}}(\omega)f_{\mu}(\bm{r},\omega)=\lambda_{\mu}(\omega)f_{\mu}(\bm{r},\omega)~{},$ (9)
where $\lambda_{\mu}(\omega)$ are the eigenvalues associated with these basis functions. The completeness and orthonormality of the basis set are given by:
$\sum_{\mu}f_{\mu}(\bm{r},\omega)f^{*}_{\mu}(\bm{r}^{\prime},\omega)=\delta(\bm{r}-\bm{r}^{\prime})~{},$ (10)
[…]
Using Eq. (10) ,we can expand the external density $\rho_{\rm ext}(\bm{r},\omega)$ in terms of the basis set:
$\rho_{\rm ext}(\bm{r},\omega)=\sum_{\mu}a_{\mu}f_{\mu}(\bm{r},\omega)~{},$ (12)
[…]
The coefficients $a_{\mu}$ are projections of the charge density onto the basis functions. It can be seen by inspection that the potential:
$\phi(\bm{r},\omega)=\sum_{\mu}\frac{4\pi a_{\mu}f_{\mu}(\bm{r},\omega)}{\lambda_{\mu}(\omega)}~{},$ (14)
solves Eq. (1). By comparing Eq. (12) and Eq. (14) with Eq. (7), we find the following mode expansion for the Green’s function:
${\rm Re}~{}g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)=4\pi\sum_{\mu}\frac{f_{\mu}(\bm{r},\omega)f^{*}_{\mu}(\bm{r}^{\prime},\omega)}{\lambda_{\mu}(\omega)}~{}.$ (15)
The poles of the Green’s function correspond to the condition $\lambda_{\mu}(\omega)=0$ and are determined by the frequencies $\omega_{\mu,s}$, where $s$ indexes different solutions of $\lambda_{\mu}(\omega)=0$. Assuming the mode function is frequency independent, $f_{\mu}(\bm{r},\omega)=f_{\mu}(\bm{r})$, we calculate the imaginary part of the Green’s function ${\rm Im}~{}g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)$ by substituting $\lambda_{\mu}(\omega)\to\lambda_{\mu}(\omega+i\eta)$ and taking the limit $\eta\to 0$. In calculating ${\rm Im}~{}g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)$, we are allowed to expand the denominator around positive (negative) poles $\omega_{\mu,s}$ ($-\omega_{\mu,s}$) as:
$\begin{split}\lambda_{\mu}(\omega+i\eta)\approx\partial_{\omega}\lambda_{\mu}(\omega)|_{\omega=\omega_{\mu,s}}\left(\omega+\mathrm{i}\eta-\omega_{\mu,s}\right)~{},\\ \lambda_{\mu}(\omega+i\eta)\approx-\partial_{\omega}\lambda_{\mu}(\omega)|_{\omega=\omega_{\mu,s}}\left(\omega+\mathrm{i}\eta+\omega_{\mu,s}\right)~{},\end{split}$ (16)
where the first equation holds around the positive pole and the second equation holds around the negative pole. We assumed that $\lambda_{\mu}(\omega)$ is an even function of $\omega$, making $\partial_{\omega}\lambda_{\mu}(\omega)$ an odd function of $\omega$. In the limit $\eta\to 0$, this yields:
$\begin{split}&{\rm Im}~{}g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)=\\ &=-4\pi^{2}\sum_{\mu,s}\frac{f_{\mu}(\bm{r})f^{*}_{\mu}(\bm{r}^{\prime})}{\partial_{\omega}\lambda_{\mu}(\omega)|_{\omega=\omega_{\mu,s}}}\left[\delta(\omega-\omega_{\mu,s})-\delta(\omega+\omega_{\mu,s})\right]~{}.\end{split}$ (17)
Being a causal response function, the Green’s function is expected to fulfill the Kramers-Kronig relations Giuliani_and_Vignale and therefore it should vanish in the limit $\omega\to\infty$. However, the term
$\begin{split}&g_{\infty}^{(0)}(\bm{r},\bm{r}^{\prime})\equiv{\rm lim}_{\omega\to\infty}g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)~{}.\\ \end{split}$ (18)
is generally finite. In the high-frequency limit, far from resonant poles, allowing us to consider $g_{\infty}^{(0)}(\bm{r},\bm{r}^{\prime})$ as a real function. However, the real and imaginary parts of the difference $g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)-g_{\infty}^{(0)}(\bm{r},\bm{r}^{\prime})$, being causal response function that goes to zero for $\omega\to\infty$, satisfy the Kramers-Kronig relations Giuliani_and_Vignale :
${\rm Re}~{}g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)=$
[…]
${\rm Im}~{}g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)=$
[…]
We now discuss the specific case of our hPP cavity. Details on the dielectric tensor in hBN are reported in Appendix A. Here, the mode functions solve the following eigenvalue equation:
$-\big{[}\epsilon_{z}(\omega)\partial^{2}_{z}+\epsilon_{x}(\omega)\nabla^{2}_{\parallel}\big{]}f_{\mu}(\bm{r},\omega)=\lambda_{\mu}(\omega)f_{\mu}(\bm{r},\omega)~{},$ (21)
with the boundary condition $f_{\mu}(\bm{r},\omega)|_{z=L_{z}/2}=f_{\mu}(\bm{r},\omega)|_{z=-L_{z}/2}=0$.
The mode functions and eigenvalues for this system are:
$f_{\bm{q}_{\parallel},n_{z}}(\bm{r})=\sqrt{\frac{2}{L_{z}S}}e^{i\bm{q}_{\parallel}\cdot\bm{r}_{\parallel}}\cos\left(\frac{n_{z}\pi z}{L_{z}}\right)~{},$ (22)
$\lambda_{\bm{q}_{\parallel},n_{z}}(\omega)=\epsilon_{z}(\omega)q_{z,n_{z}}^{2}+\epsilon_{x}(\omega)q_{\parallel}^{2}~{},$ (23)
where $\mu=\bm{q}_{\parallel},n_{z}$ indicate the modes, $\bm{q}_{\parallel}$ is the in-plane wavenumber, and , $q_{z,n_{z}}=\left[{(2n_{z}+1)\pi}/{L_{z}}\right]$, $n_{z}$ being positive integers. By means of Eq. (15) we can obtain the mode expansions of the scalar Green’s function:
$\begin{split}&g^{(0)}(\bm{r}_{\parallel},\bm{r}^{\prime}_{\parallel},z,z^{\prime},\omega)=\\ =&\frac{8\pi}{L_{z}S}\sum_{\bm{q}_{\parallel},n_{z}}\frac{e^{i\bm{q}_{\parallel}\cdot(\bm{r}_{\parallel}-\bm{r}^{\prime}_{\parallel})}\cos\left(q_{z,n_{z}}z\right)\cos\left(q_{z,n_{z}}z^{\prime}\right)}{\epsilon_{z}(\omega)q_{z,n_{z}}^{2}+\epsilon_{x}(\omega)q_{\parallel}^{2}}~{}.\end{split}$ (24)
Poles of Eq. (LABEL:eqn:green_function_hBN_expansion) are determined by $\lambda_{\bm{q}_{\parallel},n_{z}}(\omega)=0$, or, equivalently:
$\frac{\epsilon_{z}(\omega)}{\epsilon_{x}(\omega)}=-\left(\frac{q_{\parallel}L_{z}}{\pi(2n_{z}+1)}\right)^{2}~{}.$ (25)
This equation is solved for $\omega=\omega_{{q}_{\parallel},q_{z,n_{z}},s}$ where $s=1$ ($s=-1$) denotes the upper (lower) Restrahlenband. From this equation, it is clear that: i) the poles are only present in the hyperbolic region where $\epsilon_{z}(\omega)\epsilon_{x}(\omega)<0$ ; ii) the position of the poles depends only on the product $q_{\parallel}L_{z}$ rather than on the two parameters individually. Fig. 5(a) shows the cavity energies $\hbar\omega_{{q}_{\parallel},q_{z,n_{z}},s}$ associated with the poles of Eq. (LABEL:eqn:green_function_hBN_expansion) as a function of $q_{\parallel}L_{z}$ for different values of $n_{z}$.
By performing the Fourier transform of Eq. (LABEL:eqn:green_function_hBN_expansion) we obtain Eq. (LABEL:eqn:green_function_hBN_expansion_Main). Along the plane $z=0$ (and thus also $z^{\prime}=0$) the sum over $n_{z}$ can be performed analytically In the case of the scalar Green’s function, it results in
$\begin{split}&g^{(0)}(q_{\parallel},0,0,\omega)=\frac{2\pi}{q_{\parallel}\sqrt{{\epsilon}_{x}(\omega){\epsilon}_{z}(\omega)}}\tanh\left[\sqrt{\frac{{\epsilon}_{x}(\omega)}{{\epsilon}_{z}(\omega)}}\frac{L_{z}q_{\parallel}}{2}\right]~{}.\end{split}$ (26)
This expression describes the Coulomb interaction within a cavity, modified by the dielectric environment characterized by the permittivities ${\epsilon}_{x}(\omega)$ and ${\epsilon}_{z}(\omega)$, along the $x$ and $z$ directions, respectively. The presence of the mirrors at the cavity boundaries plays a critical role by effectively introducing image charges that modify the electron-electron interactions. This alteration leads to the appearance of the hyperbolic tangent term, which reflects the confinement effect of the cavity along the $z$-axis. It is also useful to recast this equation in the following form:
$\begin{split}&g^{(0)}(q_{\parallel},0,0,\omega)=\frac{2\pi L_{z}}{q_{\parallel}L_{z}\sqrt{{\epsilon}_{x}(\omega){\epsilon}_{z}(\omega)}}\tanh\left[\sqrt{\frac{{\epsilon}_{x}(\omega)}{{\epsilon}_{z}(\omega)}}\frac{L_{z}q_{\parallel}}{2}\right]~{},\end{split}$ (27)
where it is apparent that, for a fixed product $q_{\parallel}L_{z}$, the Green’s function scales as $L_{z}$. In the limit $L_{z}\to\infty$, the Green’s function correctly reduces to the free-space Coulomb potential renormalized by the frequency-dependent dielectric properties of hBN:
$\begin{split}g_{L_{z}\to\infty}^{(0)}({q}_{\parallel},0,0,\omega)=\frac{2\pi}{q_{\parallel}\sqrt{\epsilon_{x}(\omega)\epsilon_{z}(\omega)}}~{}.\end{split}$ (28)
In the limit $\omega\to\infty$, the scalar Green’s function reads
$\begin{split}&g_{\infty}^{(0)}(q_{\parallel},0,0)=\frac{2\pi}{q_{\parallel}\sqrt{{\epsilon}_{x,\infty}{\epsilon}_{z,\infty}}}\tanh\left[\sqrt{\frac{{\epsilon}_{x,\infty}}{{\epsilon}_{z,\infty}}}\frac{L_{z}q_{\parallel}}{2}\right]~{}.\end{split}$ (29)
[…]
Figure 5: (Color online) Panel (a) shows the cavity frequencies $\hbar\omega_{q_{\parallel},n_{z},s}$ (in ${\rm eV}$) as a function of $q_{\parallel}L_{z}$ for different $n_{z}$. Each of the two regions corresponding to the Reststrahlen bands is enclosed between two black dashed lines. Panel (b) displays the bare cavity vacuum fluctuations $e^{2}\phi^{2}_{{q}_{\parallel},n_{z},s}(0)$ (normalized by $L_{z}$ and expressed in ${\rm nm\times eV}^{2}$) as a function of $q_{\parallel}L_{z}$ for different $n_{z}$. Panel (c) illustrates $e^{2}\phi^{2}_{{q}_{\parallel},n_{z},s}(0)$ (expressed in ${\rm nm}^{2}\times{\rm eV}^{2}$) as a function of $L_{z}$ (in nm) for different $n_{z}$. In all panels, the upper (lower) Reststrahlen band is shown in blue (red). The thickness of the lines decreases as the quantum number $n_{z}$ increases.
III.1 A quantization scheme for sub-wavelength cavities
Here, we derive a quantization scheme to associate the Green’s function in Eq. (LABEL:eqn:green_function_hBN_expansion_Main) with the corresponding quantum mechanical modes associated with the resonances $\hbar\omega_{{q}_{\parallel},n_{z},s}$. Before addressing the specific case of hBN sub-wavelength cavities, we consider a generic scenario with a resonance $\hbar\omega_{\mu,s}$, where $\mu$ is a generic index encompassing different quantum numbers (e.g. $\mu=\bm{q}_{\parallel},n_{z}$ in the hBN case) and $s$ denotes different solutions with the same mode function.
The cavity Hamiltonian, associated with these poles, is given by:
^$\hat{\mathcal{H}}_{\rm cav}=\sum_{\mu,s}\hbar\omega_{\mu,s}\hat{a}^{\dagger}_{\mu,s}\hat{a}_{\mu,s}~{}.$ (30)
We assume that the scalar potential can be expanded in terms of the modes $\hat{a}^{\dagger}_{\mu,s}$ and $\hat{a}_{\mu,s}$ as:
$\hat{\phi}(\bm{r})=\sum_{\mu,s}\frac{1}{\sqrt{N_{\mu,s}}}\left(f_{\mu}(\bm{r})\hat{a}_{\mu,s}+f^{*}_{\mu}(\bm{r})\hat{a}^{\dagger}_{\mu,s}\right)~{},$ (31)
where $N_{\mu,s}$ are normalization constants that will be determined later. Eq. (31) allows us to compute the quantum mechanical fluctuations of the field:
$\tensor{[}_{\rm cav}]{\langle}{}0|\hat{\phi}(\bm{r})\hat{\phi}(\bm{r}^{\prime})|0\tensor{[}]{\rangle}{{}_{\rm cav}}$ $=\sum_{\mu,s}\frac{f_{\mu}(\bm{r})f^{*}_{\mu}(\bm{r}^{\prime})}{N_{\mu,s}}~{},$ (32)
where $|0\tensor{[}]{\rangle}{{}_{\rm cav}}$ is the vacuum of the cavity Hamiltonian ^$\hat{\mathcal{H}}_{\rm cav}$. Assuming the mode expansion in Eq. (31) and the cavity Hamiltonian ^$\hat{\mathcal{H}}_{\rm cav}$ in Eq. (30), we can use the Kubo formula Giuliani_and_Vignale to calculate the scalar Green function associated with the modes, $g_{\rm q}^{(0)}(\bm{r},\bm{r}^{\prime},\tau)$:
$\begin{split}g_{\rm q}^{(0)}(\bm{r},\bm{r}^{\prime},\tau)\equiv-\frac{\mathrm{i}}{\hbar}\Theta(\tau)\tensor{[}_{\rm cav}]{\langle}{}0|\big{[}\hat{\phi}^{\rm I}(\bm{r},\tau),\hat{\phi}(\bm{r}^{\prime})\big{]}|0\tensor{[}]{\rangle}{{}_{\rm cav}}~{},\end{split}$ (33)
where $\Theta(\tau)$ is the Heaviside function, and $[\hat{X},\hat{Y}]$ denotes the commutator between the operators $\hat{X}$ and $\hat{Y}$. The term $\hat{\phi}^{\rm I}(\bm{r},t)$ refers to the electric potential in the interaction picture, specifically defined as $\hat{\phi}^{\rm I}(\bm{r},t)\equiv U_{0}^{\dagger}(t)\hat{\phi}(\bm{r})U_{0}(t)$, where $U_{0}(t)\equiv\exp(-\mathrm{i}\hat{\mathcal{H}}_{\rm cav}t/\hbar)$.
By taking the time Fourier transform of the previous equation,
[…]
and utilizing the Hermitian property, $f_{\mu}(\bm{r})f^{*}_{\mu}(\bm{r}^{\prime})=f_{\mu}(\bm{r}^{\prime})f^{*}_{\mu}(\bm{r})$, we obtain:
$\begin{split}&g_{\rm q}^{(0)}(\bm{r},\bm{r}^{\prime},\omega)=\sum_{\mu,s}\frac{f_{\mu}(\bm{r})f^{*}_{\mu}(\bm{r}^{\prime})}{{N}_{\mu,s}}\times\\ &\times\left(\frac{1}{\hbar\omega-\hbar\omega_{\mu,s}+\mathrm{i}\hbar\eta}-\frac{1}{\hbar\omega+\hbar\omega_{\mu,s}+\mathrm{i}\hbar\eta}\right)~{}.\end{split}$ (35)
This expression reveals how the quantum Green’s function is linked to the vacuum virtual excitation of the quantum field $\hat{\phi}(\bm{r}^{\prime})$. In fact,
$\tensor{[}_{\rm cav}]{\langle}{}0|\hat{\phi}(\bm{r}^{\prime})|\mu,s\tensor{[}]{\rangle}{{}_{\rm cav}}=f_{\mu}(\bm{r})/\sqrt{{N}_{\mu,s}}$ (36)
were $|\mu,s\tensor{[}]{\rangle}{{}_{\rm cav}}=a_{\mu,s}^{\dagger}|0\tensor{[}]{\rangle}{{}_{\rm cav}}$ are single-photon states.
This can interpreted within the general framework of the fluctuation-dissipation theorem Giuliani_and_Vignale , which connects ${\rm Im}\,g_{\rm q}^{(0)}(\bm{r},\bm{r}^{\prime},\omega)$ to the spectrum of the fluctuations $S^{(0)}_{\rm cav}(\bm{r},\bm{r}^{\prime},\omega)$,
$\begin{split}&{\rm Im}~{}g_{\rm q}^{(0)}(\bm{r},\bm{r}^{\prime},\omega)=-\frac{\pi}{\hbar}S^{(0)}_{\rm cav}(\bm{r},\bm{r}^{\prime},\omega)~{}.\end{split}$ (37)
where
$\begin{split}&S^{(0)}_{\rm cav}(\bm{r},\bm{r}^{\prime},\omega)\equiv\frac{1}{2\pi}\int_{-\infty}^{\infty}d\tau e^{\mathrm{i}\omega\tau}\tensor{[}_{\rm cav}]{\langle}{}0|\hat{\phi}^{\rm I}(\bm{r},\tau)\hat{\phi}(\bm{r}^{\prime})|0\tensor{[}]{\rangle}{{}_{\rm cav}}~{}.\end{split}$ (38)
Note that $S^{(0)}_{\rm cav}(\bm{r},\bm{r}^{\prime},\omega)$ encodes the fluctuation of the field in Eq. (32) as $\int_{-\infty}^{\infty}d\omega S^{(0)}_{\rm cav}(\bm{r},\bm{r}^{\prime},\omega)=\tensor{[}_{\rm cav}]{\langle}{}0|\hat{\phi}(\bm{r})\hat{\phi}(\bm{r}^{\prime})|0\tensor{[}]{\rangle}{{}_{\rm cav}}$.
By taking the real and imaginary part of Eq. (LABEL:eq:green_q), and the limit $\eta\to 0$, we have:
$\begin{split}&{\rm Re}\,g_{\rm q}^{(0)}(\bm{r},\bm{r}^{\prime},\omega)=\sum_{\mu}\frac{f_{\mu}(\bm{r})f^{*}_{\mu}(\bm{r}^{\prime})}{N_{\mu}}\left[\frac{2\hbar\omega_{\mu}}{(\hbar\omega)^{2}-(\hbar\omega_{\mu})^{2}}\right]~{},\end{split}$ (39)
$\begin{split}&{\rm Im}\,g_{\rm q}^{(0)}(\bm{r},\bm{r}^{\prime},\omega)=\\ &=-\pi\sum_{\mu}\frac{f_{\mu}(\bm{r})f^{*}_{\mu}(\bm{r}^{\prime})}{\hbar N_{\mu}}\left[\delta(\omega-\omega_{\mu})-\delta(\omega+\omega_{\mu})\right]~{},\end{split}$ (40)
By comparing Eq. (LABEL:eq:GM_im) with Eq. (LABEL:eqn:green_function_IM), by choosing
$\begin{split}N_{\mu,s}=\frac{1}{4\pi\hbar}\partial_{\omega}\lambda_{\mu}(\omega)|_{\omega=\omega_{\mu,s}}~{},\end{split}$ (41)
we can identify the imaginary part of the Green’s function ${\rm Im}~{}g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)$ with the imaginary part of the mode Green’s function ${\rm Im}~{}g_{\rm q}^{(0)}(\bm{r},\bm{r}^{\prime},\omega)$:
$\begin{split}{\rm Im}\,g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)={\rm Im}\,g_{\rm q}^{(0)}(\bm{r},\bm{r}^{\prime},\omega)~{}.\end{split}$ (42)
As $g_{\rm q}^{(0)}(\bm{r},\bm{r}^{\prime},\omega)$ fulfills the Kramers-Kronig relations Giuliani_and_Vignale :
[…]
using Eq. (42), we find:
$\begin{split}&g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)=g_{\rm q}^{(0)}(\bm{r},\bm{r}^{\prime},\omega)+g_{\infty}^{(0)}(\bm{r},\bm{r}^{\prime})~{}.\end{split}$ (45)
The total Green’s function can be decomposed as a sum of the dynamical quantum Green’s function $g_{\rm q}^{(0)}(\bm{r},\bm{r}^{\prime},\omega)$ and the static potential $g_{\infty}^{(0)}(\bm{r},\bm{r}^{\prime})$.
We now specialize these results for the case of hBN. The normalization factors $N_{q_{\parallel},n_{z},s}$ can be expressed explicitly in terms of the dielectric function of hBN as follows:
$\begin{split}&N_{q_{\parallel},n_{z},s}=\\ &=\frac{1}{4\pi\hbar}\left[\partial_{\omega}\epsilon_{z}(\omega)q_{z,n_{z}}^{2}+\partial_{\omega}\epsilon_{x}(\omega)q_{\parallel}^{2}\right]_{\omega=\omega_{q_{\parallel},n_{z},s}}~{}.\end{split}$ (46)
By applying Eq. (31) to the hBN cavity, we obtain Eq. (3) while the explicit expression of the normalization factors $N_{q_{\parallel},n_{z},s}$ is given in Eq. (LABEL:eq:N_q). The Green’s function can be expressed in terms of the coefficients $\phi_{q_{\parallel},n_{z},s}(z)$ as:
$\begin{split}&g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)=\sum_{n_{z},s}\phi_{{q}_{\parallel},n_{z},s}(z)\phi_{{q}_{\parallel},n_{z},s}(z^{\prime})\times\\ &\times\left(\frac{1}{\hbar\omega-\hbar\omega_{q_{\parallel},n_{z},s}+\mathrm{i}\hbar\eta}-\frac{1}{\hbar\omega+\hbar\omega_{q_{\parallel},n_{z},s}+\mathrm{i}\hbar\eta}\right)~{}.\\ \end{split}$
Additionally, the vacuum fluctuations of the field are related to the coefficients $\phi_{q_{\parallel},n_{z},s}(z)$:
$\tensor{[}_{\rm cav}]{\langle}{}0|\hat{\phi}^{2}(\bm{r}_{\parallel},0)|0\tensor{[}]{\rangle}{{}_{\rm cav}}$ $=\sum_{\bm{q}_{\parallel},n_{z},s}\frac{1}{S}\phi^{2}_{{q}_{\parallel},n_{z},s}({0})~{}.$ (47)
Hence, the coefficients $\phi_{q_{\parallel},n_{z},s}(z)$ can be interpreted as a measure of the density of vacuum fluctuations. Moreover, these coefficients also appear in the Green’s function, as shown in Eq. (LABEL:eqn:green_function_hBN_expansion_quantum), indicating that the intensity of a scalar field generated by placing a test charge depends on the vacuum fluctuations of the field in the absence of any charge, represented by $\phi_{q_{\parallel},n_{z},s}(z)$. It thus is useful to express $\phi_{q_{\parallel},n_{z},s}(z)$ in the following form:
$\phi^{2}_{{q}_{\parallel},n_{z},s}({0})=\frac{8\pi\hbar L_{z}}{\left[\partial_{\omega}\epsilon_{z}(\omega)[(2n_{z}+1)\pi]^{2}+\partial_{\omega}\epsilon_{x}(\omega)(q_{\parallel}L_{z})^{2}\right]_{\omega=\omega_{q_{\parallel},n_{z},s}}}~{},$ (48)
where we used $q_{z,n_{z}}L_{z}=(2n_{z}+1)\pi$. Note that $\phi^{2}_{q_{\parallel},n_{z},s}(0)/L_{z}$ depends only on the product $q_{\parallel}L_{z}$ and not on these variables individually. Fig. 5(b,c) shows the vacuum fluctuation coefficients $\phi_{q_{\parallel},n_{z},s}(z)$ as a function of $q_{\parallel}L_{z}$ in panel (b) and as a function of $L_{z}$ in panel (c). As seen in panel (b), the maximum of $\phi_{q_{\parallel},n_{z},s}(z)$ occurs at $q_{\parallel}L_{z}=0$ for the lower Reststrahlen band, while for the upper Reststrahlen band, the maximum is reached when $q_{\parallel}L_{z}$ is on the order of unity. Given that the vacuum fluctuation coefficients quantify the intensity of the field’s response to external perturbations, this implies that the field response is maximized under the same conditions: at $q_{\parallel}L_{z}=0$ for the lower Reststrahlen band and at $q_{\parallel}L_{z}\sim 1$ for the upper band.
Finally, in the HPP cavity under study, Eq. (LABEL:eqn:green_function_hBN_expansion_R) reads:
$\begin{split}&g^{(0)}(q_{\parallel},z,z^{\prime},\omega)=g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)+g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})~{},\end{split}$ (49)
where $g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})$ represents the high-frequency limit of the Green’s function (also given in Eq. (LABEL:eqn:green_function_hBN_classical_inf) where the sum over $n_{z}$ is carried out):
$\begin{split}&g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})=\frac{8\pi}{L_{z}}\sum_{n_{z}=0}^{\infty}\frac{\cos\left(q_{z,n_{z}}z\right)\cos\left(q_{z,n_{z}}z^{\prime}\right)}{{\epsilon}_{z,\infty}q_{z,n_{z}}^{2}+{\epsilon}_{x,\infty}q_{\parallel}^{2}}~{}.\end{split}$ (50)
This means that the quantum expansion of the scalar potential (3) reproduces the correct Green’s function $g^{(0)}(q_{\parallel},z,z^{\prime},\omega)$, in Eq. (LABEL:eqn:green_function_hBN_expansion), provided the static potential $g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})$ is added to the Hamiltonian as the standard static Coulomb interaction:
^$\hat{\mathcal{H}}_{\rm C}=\frac{1}{2S}\sum_{\bm{q}_{\parallel}}\hat{n}_{-\bm{q}_{\parallel}}e^{2}g_{\infty}^{(0)}(q_{\parallel},0,0)\hat{n}_{\bm{q}_{\parallel}}~{},$ (51)
where $\hat{n}_{\bm{q}_{\parallel}}$ is the wavevector-resolved density operator
[…]
IV Details on “Theory of Landau levels in Graphene and their coupling to a subwavelength cavity”
[…]
Figure 6: (Color online) Panel (a) shows the energy differences $\varepsilon_{\lambda,n}-\varepsilon_{\lambda^{\prime},n^{\prime}}$ associated with the LL transitions as a function of the ratio $\hbar v_{\rm F}/\ell_{B}$. Panel (b) displays the squared modulus of the density matrix element, $\ell^{2}_{B}\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}(q_{\parallel})$, as a function of $q_{\parallel}\ell_{B}$. In both panels, each transition is labeled in the legend by the quantum numbers $(\lambda n,\lambda^{\prime}n^{\prime})$, and solid (dashed) lines indicate dipole-allowed (dipole-forbidden) transitions.
In this Section, we review the LLs in graphene under a uniform magnetic field $B$ and their coupling to a subwavelength cavity. We derive the Hamiltonian governing LLs, emphasizing the eigenstates and associated energies $\varepsilon_{\lambda,n}$, which depend on the magnetic field strength via the magnetic length $\ell_{B}$. Subsequently, we use eigenstates and eigenenergies to calculate the density–density response function $\chi^{(0)}_{nn}(q_{\parallel},\omega)$. These findings will be useful later to calculate the polaritonic specturm.
The Hamiltonian describing graphene electrons in presence of a uniform magnetic field $B$ reads castroneto_rmp_2009 ; Goerbig11 :
^$\hat{\mathcal{H}}_{\rm LLs}=v_{\rm F}\hat{\bm{\pi}}_{\parallel}\cdot\hat{\bm{\sigma}}_{\parallel}~{},$ (53)
where $v_{\rm F}$ is the Fermi velocity, $\hat{\bm{\sigma}}_{\parallel}$ are the Pauli-operators acting on the Valley degrees of freedom and the kinetic momentum operators $\hat{\pi}_{x},\hat{\pi}_{y}$ are
$\begin{split}\hat{\pi}_{x}&=\hat{p}_{x}-\frac{eB}{c}\hat{y}~{},\\ \hat{\pi}_{y}&=\hat{p}_{y}~{}.\\ \end{split}$ (54)
Here, we have chosen the Landau gauge where the vector potential reads $\bm{A}(\bm{r})=(A_{x}(\bm{r}),A_{y}(\bm{r}))=(-By,0)$. With this gauge choice, $\hat{x}$ does not appear in the Hamiltonian, which allows us to use the plane wave states |$\ket{k_{x}}$, such that $\braket{x}{k_{x}}=e^{-\mathrm{i}k_{x}x}/\sqrt{L_{x}}$ and $\hat{p}_{x}\ket{k_{x}}=\hbar k_{x}\ket{k_{x}}$. In terms of these operators, the LLs Hamiltonian ^$\hat{\mathcal{H}}_{\rm LLs}$ reads:
^$\begin{split}\hat{\mathcal{H}}_{\rm LLs}&=v_{\rm F}\left[\hbar\left(k_{x}-\hat{y}/{{\ell}^{2}_{B}}\right)\hat{\sigma}_{x}+\hat{\pi}_{y}\hat{\sigma}_{y}\right]~{},\end{split}$ (55)
where ${\ell}_{B}=\sqrt{\left({\hbar c}/{eB}\right)}$ is the magnetic length. By defining the displaced position $\hat{{y}}_{k_{x}}$
$\begin{split}\hat{{y}}_{k_{x}}&=\hat{y}-k_{x}\ell_{B}^{2}~{},\end{split}$ (56)
we can define the ladder operators $\hat{a}_{k_{x}},\hat{a}_{k_{x}}^{\dagger}$ associated with $\hat{{y}}_{k_{x}}$ such that
$\begin{split}\hat{{y}}_{k_{x}}&=\frac{\ell_{B}}{\sqrt{2}}\left(\hat{a}_{k_{x}}+\hat{a}_{k_{x}}^{\dagger}\right)~{},\\ \hat{p}_{x}&=\frac{\hbar\mathrm{i}}{\sqrt{2}\ell_{B}}\left(\hat{a}_{k_{x}}^{\dagger}-\hat{a}_{k_{x}}\right)~{}.\end{split}$ (57)
In terms of these ladder operators the LLs Hamiltonian ^$\hat{\mathcal{H}}_{\rm LLs}$ reads:
^$\begin{split}\hat{\mathcal{H}}_{\rm LLs}&=-\frac{\hbar v_{\rm F}\sqrt{2}}{\ell_{B}}\begin{pmatrix}0&\hat{a}_{k_{x}}\\ \hat{a}^{\dagger}_{k_{x}}&0\end{pmatrix}\end{split}~{}.$ (58)
It can be seen by inspection that the eigenvector of the LLs Hamiltonian ^$\hat{\mathcal{H}}_{\rm LLs}$ is where |$\ket{\lambda,k_{x},n}$ are the following eigenstates
[…]
with $\lambda=\pm$ and
|$\ket{n_{k_{x}}}=\frac{(\hat{a}_{k_{x}}^{\dagger})^{n}}{\sqrt{n!}}\ket{0_{k_{x}}}$ (60)
and |$\ket{0_{k_{x}}}$ is the $k_{x}$-dependent vacuum, such that $\hat{a}_{k_{x}}\ket{0_{k_{x}}}=0$, and
$C^{\pm}_{n}=\sqrt{\frac{1\pm\delta_{0n}}{2}}~{}.$ (61)
The associated energies are $\varepsilon_{\lambda,n}\equiv\lambda(\hbar v_{\rm F}/\ell_{B})\sqrt{2n}$. Note that these energies do not depend on the quantum number $k_{x}$, which reflects the degeneracy of the Landau levels in the $x$-direction. Notice that $n=0$ corresponds to only one state, without the degeneracy $\lambda=\pm$. Conventionally, we associate the quantum number $\lambda=-$ with $n=0$. Each LL has a degeneracy of
$N_{\text{L}}=\frac{gS}{2\pi\ell_{B}^{2}}~{},$ (62)
where $g=g_{\text{s}}g_{\text{v}}$ takes into account the spin and valley degeneracies.
Realistic values of the parameters for the graphene Hamiltonian in a magnetic field castroneto_rmp_2009 ; Goerbig11 are provided in Table 1, while the ranges of the magnetic field $B$, magnetic length $\ell_{B}$, and cyclotron frequency $\hbar v_{\rm F}/\ell_{B}$ used in the paper are summarized in Table 2.
$v_{\rm F}$ $10^{6}~{}{\rm m/s}$
$\ell_{B}$ $25.6/\sqrt{B[{\rm T}]}~{}{\rm nm}$
$\hbar v_{\rm F}/\ell_{B}$ $25.7\sqrt{B[{\rm T}]}~{}{\rm meV}$
Table 1: Realistic values of the parameters for the graphene Hamiltonian coupled with a magnetic field $B$ castroneto_rmp_2009 ; Goerbig11 .
$B$ $5.4-49.1{\rm T}$
$\ell_{B}$ $11.0-3.7~{}{\rm nm}$
$\hbar v_{\rm F}/\ell_{B}$ $0.04-0.18~{}{\rm eV}$
Table 2: Values of parameters relative to LLs spanned in this article.
Fig. 6(a) shows the energy differences $\varepsilon_{\lambda,n}-\varepsilon_{\lambda^{\prime},n^{\prime}}$ associated with the LL transitions as a function of the ratio $\hbar v_{\rm F}/\ell_{B}$. Since $\ell_{B}$ depends on the magnetic field, it is experimentally feasible to adjust $\ell_{B}$ by varying the field strength. This allows for precise control over the LL transition energies, enabling them to be tuned into resonance with the cavity modes.
In Appedix B we calculate the bare density-density response function, $\chi^{(0)}_{nn}(q_{\parallel},\omega)$. This is relevant as the coupling with the cavity is through the density operator as in Eq. (4). The density–density response function reads roldan2009 ; roldan2010
$\chi^{(0)}_{nn}(q_{\parallel},\omega)=\sum_{n,\lambda,n^{\prime},\lambda^{\prime}}\frac{(f_{n\lambda}-f_{n^{\prime}\lambda^{\prime}})\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})}{\hbar\omega+\varepsilon_{\lambda,n}-\varepsilon_{\lambda^{\prime},n^{\prime}}+\mathrm{i}\eta}~{}.$ (63)
where $\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})$ represents the squared modulus of the density matrix element,
$\begin{split}&\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})\equiv\sum_{k^{\prime}_{x},k_{x}}\lvert\braket{k_{x}^{\prime},n^{\prime},\lambda^{\prime}}{\hat{n}_{\bm{q}_{\parallel}}}{k_{x},n,\lambda}\rvert^{2}~{},\\ &=\frac{g}{2\pi\ell_{B}^{2}}e^{-q_{\parallel}^{2}\ell_{B}^{2}/2}|\mathcal{F}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})|^{2}~{},\end{split}$ (64)
where $\mathcal{F}_{n^{\prime},\lambda^{\prime},n,\lambda}(q_{\parallel})$ is defined in Appendix B. Fig. 6 (b) shows the squared modulus of the density matrix element, $\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}(q_{\parallel})$, as a function of $q_{\parallel}\ell_{B}$. The maximum of $\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}(q_{\parallel})$ is reached when $q_{\parallel}\ell_{B}$ is on the order of unity for all the considered LL transitions.
At zero temperature, $T=0$, the density–density response function is
$\chi^{(0)}_{nn}(q_{\parallel},\omega)=\sum_{n,\lambda\in\mathcal{N}_{\rm occ}}\sum_{n^{\prime},\lambda^{\prime}\in\mathcal{N}_{\rm unocc}}\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})\left(\frac{1}{\hbar\omega+\varepsilon_{\lambda,n}-\varepsilon_{\lambda^{\prime},n^{\prime}}+\mathrm{i}\eta}-\frac{1}{{\hbar\omega+\varepsilon_{\lambda^{\prime},n^{\prime}}-\varepsilon_{\lambda,n}+\mathrm{i}\eta}}\right)~{},$ (65)
where the sums run over occupied ($\mathcal{N}_{\rm occ}$) and unoccupied ($\mathcal{N}_{\rm unocc}$) LLs, respectively.
V Details on “The polaritonic spectrum and the vacuum Rabi frequency”
In this Section, we derive and diagonalize the Hopfield Hamiltonian to describe the polaritonic spectrum of our system. The Hamiltonian includes contributions from LLs, Coulomb interactions, HPP cavity modes, and light-matter coupling. By introducing bright modes, the LL Hamiltonian is expressed in terms of bosonic operators, and the interaction with cavity modes is described using Rabi frequencies, defined by vacuum fluctuation coefficients and by matrix elements of the density operator.
V.1 Diagonalizing the Hopfield Hamiltonian
We now proceed to derive and diagonalize the Hopfield polaritonic Hamiltonian hopfield1958 ; Kavokin11 for our system, which is given by:
^=^$\hat{\mathcal{H}}=\hat{\mathcal{H}}_{\rm LLs}+\hat{\mathcal{H}}_{\rm C}+\hat{\mathcal{H}}_{\rm cav}+\hat{\mathcal{H}}_{\rm int}~{},$ (66)
where ^$\hat{\mathcal{H}}_{\rm LLs}$ represents the Hamiltonian for the LLs, ^$\hat{\mathcal{H}}_{\rm C}$ describes the Coulomb interactions among the electrons, ^$\hat{\mathcal{H}}_{\rm cav}$ is the cavity Hamiltonian (see Eq. (30)), and ^$\hat{\mathcal{H}}_{\rm int}$ accounts for the interaction between the LLs and the cavity mode (see Eq. (4)). LL states with energy $\varepsilon_{\lambda,n}$ are created (destroyed) by the fermionic operators $\hat{c}^{\dagger}_{k_{x},\lambda,n}$ ($\hat{c}_{k_{x},\lambda,n}$). Thus the matter LL Hamiltonian reads
^$\hat{\mathcal{H}}_{\rm LLs}=\sum_{k_{x},\lambda,n}\varepsilon_{\lambda,n}\hat{c}^{\dagger}_{k_{x},\lambda,n}\hat{c}_{k_{x},\lambda,n}~{}.$ (67)
For convenience, we decompose this Hamiltonian into contributions from electrons and holes, such that ^$\hat{\mathcal{H}}_{\rm LLs}=\hat{\mathcal{H}}^{(+)}_{\rm LLs}+\hat{\mathcal{H}}^{(-)}_{\rm LLs}$ where
^$\hat{\mathcal{H}}^{(+)}_{\rm LLs}=\sum_{n^{\prime},\lambda^{\prime}\in\mathcal{N}_{\rm unocc}}\sum_{k_{x}^{\prime}}\varepsilon_{\lambda^{\prime},n^{\prime}}\hat{c}^{\dagger}_{k_{x}^{\prime},\lambda^{\prime},n^{\prime}}\hat{c}_{k_{x}^{\prime},\lambda^{\prime},n^{\prime}}~{},$ (68)
and
^$\hat{\mathcal{H}}^{(-)}_{\rm LLs}=\sum_{n,\lambda\in\mathcal{N}_{\rm occ}}\sum_{k_{x}}\varepsilon_{\lambda,n}\hat{c}^{\dagger}_{k_{x},\lambda,n}\hat{c}_{k_{x},\lambda,n}~{}.$ (69)
We now introduce the bright modes $\hat{b}^{\dagger}_{n,\lambda,n,\lambda}(q_{x})$ defined as
$\begin{split}\hat{b}^{\dagger}_{n,\lambda,n^{\prime},\lambda^{\prime}}(q_{x})&\equiv\frac{1}{\sqrt{N_{\rm L}}}\sum_{k_{x}}\hat{c}^{\dagger}_{k_{x}-q_{x},\lambda^{\prime},n^{\prime}}\hat{c}_{k_{x},\lambda,n}~{},\\ \end{split}$ (70)
where ${n^{\prime},\lambda^{\prime}\in\mathcal{N}_{\rm unocc}}$ and ${n,\lambda\in\mathcal{N}_{\rm occ}}$. Bright modes are approximatively bosonic in the dilute limit hagenmuller2010 . Indeed, the commutator reads
$\begin{split}&[\hat{b}_{n^{\prime\prime},\lambda^{\prime\prime},n^{\prime\prime\prime},\lambda^{\prime\prime\prime}}(q_{x}^{\prime}),\hat{b}^{\dagger}_{n,\lambda,n^{\prime},\lambda^{\prime}}(q_{x})]=\\ &=\frac{1}{N_{\rm L}}\sum_{k_{x}^{\prime}}\big{[}\delta_{\lambda^{\prime\prime\prime},\lambda^{\prime}}\delta_{n^{\prime\prime\prime},n^{\prime}}\hat{c}^{\dagger}_{k_{x}^{\prime},\lambda^{\prime\prime},n^{\prime\prime}}\hat{c}_{k_{x}^{\prime}-q_{x}^{\prime}+q_{x},\lambda,n}+\\ &-\delta_{\lambda,\lambda^{\prime\prime}}\delta_{n,n^{\prime\prime}}\hat{c}^{\dagger}_{k_{x}^{\prime}-q_{x},\lambda^{\prime},n^{\prime}}\hat{c}_{k_{x}^{\prime}-q_{x}^{\prime},\lambda^{\prime\prime\prime},n^{\prime\prime\prime}}\big{]}~{}.\end{split}$ (71)
We now approximate the remaining operators by their expectation values over the Fermi sea, $|{\rm FS}\tensor{[}]{\rangle}{}$, which represents the ground state of the system at zero temperature ($T=0$). Specifically, we consider the expectation value of the product of creation and annihilation operators as follows:
$\begin{split}&\hat{c}^{\dagger}_{k_{x}^{\prime},\lambda^{\prime\prime},n^{\prime\prime}}\hat{c}_{k_{x}^{\prime}-q_{x}^{\prime}+q_{x},\lambda,n}\approx\\ &\approx\tensor{[}]{\langle}{}{\rm FS}|\hat{c}^{\dagger}_{k_{x}^{\prime},\lambda^{\prime\prime},n^{\prime\prime}}\hat{c}_{k_{x}^{\prime}-q_{x}^{\prime}+q_{x},\lambda,n}|{\rm FS}\tensor{[}]{\rangle}{}=\delta_{q_{x}^{\prime},q_{x}}\delta_{\lambda^{\prime\prime},\lambda}\delta_{n^{\prime\prime},n}~{},\end{split}$ (72)
since $n,\lambda,n^{\prime\prime},\lambda^{\prime\prime}\in\mathcal{N}_{\rm occ}$ at $T=0$. Similarly, for transitions between unoccupied states, we have:
$\begin{split}&\hat{c}^{\dagger}_{k_{x}^{\prime}-q_{x},\lambda^{\prime},n^{\prime}}\hat{c}_{k_{x}^{\prime}-q_{x}^{\prime},\lambda^{\prime\prime\prime},n^{\prime\prime\prime}}\approx\\ &\approx\tensor{[}]{\langle}{}{\rm FS}|\hat{c}^{\dagger}_{k_{x}^{\prime}-q_{x},\lambda^{\prime},n^{\prime}}\hat{c}_{k_{x}^{\prime}+q_{x}^{\prime},\lambda^{\prime\prime\prime},n^{\prime\prime\prime}}|{\rm FS}\tensor{[}]{\rangle}{}=0~{},\end{split}$ (73)
for $n^{\prime},\lambda^{\prime},n^{\prime\prime\prime},\lambda^{\prime\prime\prime}\in\mathcal{N}_{\rm unocc}$ at $T=0$. Additionally, noting that $\sum_{k_{x}^{\prime}}/N_{\rm L}=1$, the commutators behave as bosonic:
$\begin{split}&[\hat{b}_{n^{\prime\prime},\lambda^{\prime\prime},n^{\prime\prime\prime},\lambda^{\prime\prime\prime}}(q_{x}^{\prime}),\hat{b}^{\dagger}_{n,\lambda,n^{\prime},\lambda^{\prime}}(q_{x})]=\\ &=\delta_{\lambda^{\prime\prime\prime},\lambda^{\prime}}\delta_{n^{\prime\prime\prime},n^{\prime}}\delta_{\lambda^{\prime\prime},\lambda}\delta_{n^{\prime},n}\delta_{q_{x}^{\prime},q_{x}}~{}.\end{split}$ (74)
Thus, the LLs Hamiltonian simplifies to:
^$\begin{split}\hat{\mathcal{H}}_{\rm LLs}\approx&\sum_{n^{\prime},\lambda^{\prime}\in\mathcal{N}_{\rm unocc}}\sum_{n,\lambda\in\mathcal{N}_{\rm occ}}\sum_{q_{x}}\big{(}\varepsilon_{\lambda,n}-\varepsilon_{\lambda^{\prime},n^{\prime}}\big{)}\\ &\times\hat{b}^{\dagger}_{n,\lambda,n^{\prime},\lambda^{\prime}}(q_{x})\hat{b}_{n,\lambda,n^{\prime},\lambda^{\prime}}(q_{x})~{},\end{split}$ (75)
which is quadratic and diagonal in terms of the bright modes. By inserting the mode expansion in Eq. (3), the light-matter interaction Hamiltonian ^$\hat{\mathcal{H}}_{\rm int}$ in Eq. (4) can be recasted as
^$\hat{\mathcal{H}}_{\rm int}=-e\sum_{\bm{q}_{\parallel},n_{z},s}\frac{\phi_{\bm{q}_{\parallel},n_{z},s}(0)}{\sqrt{S}}\left[\hat{a}^{\dagger}_{\bm{q}_{\parallel},n_{z},s}\hat{n}_{\bm{q}_{\parallel}}+\hat{a}_{\bm{q}_{\parallel},n_{z}}\hat{n}_{-\bm{q}_{\parallel}}\right]~{},$ (76)
In the density operator $\hat{n}_{\bm{q}_{\parallel}}$, we retain only the terms corresponding to the creation of electron-hole pairs, neglecting other processes:
$\begin{split}\hat{n}_{\bm{q}_{\parallel}}&\approx\sum_{n^{\prime},\lambda^{\prime}\in\mathcal{N}_{\rm unocc}}\sum_{n,\lambda\in\mathcal{N}_{\rm occ}}\sum_{k_{x},k_{x}^{\prime}}\times\\ &\times\big{(}\braket{k_{x}^{\prime},n^{\prime},\lambda^{\prime}}{\hat{n}_{\bm{q}_{\parallel}}}{k_{x},n,\lambda}\hat{c}^{\dagger}_{k_{x}^{\prime},n^{\prime},\lambda^{\prime}}\hat{c}_{k_{x},n,\lambda}+\\ &+\braket{k_{x},n,\lambda}{\hat{n}_{\bm{q}_{\parallel}}}{k_{x}^{\prime},n^{\prime},\lambda^{\prime}}\hat{c}^{\dagger}_{k_{x},n,\lambda}\hat{c}_{k_{x}^{\prime},n^{\prime},\lambda^{\prime}}\big{)}~{}.\end{split}$ (77)
Next, we incorporate the explicit form of the matrix element of the density operator, as obtained in Appendix B:
$\begin{split}&\braket{k_{x}^{\prime},n^{\prime},\lambda^{\prime}}{\hat{n}_{\bm{q}_{\parallel}}}{k_{x},n,\lambda}=\\ &=e^{-q_{\parallel}^{2}\ell_{B}^{2}/4}\delta_{k_{x}^{\prime},k_{x}-q_{x}}\mathcal{F}_{n^{\prime},\lambda^{\prime},n,\lambda}(\bm{q}_{\parallel})~{}.\end{split}$ (78)
Substituting this result into the expression for $\hat{n}_{\bm{q}_{\parallel}}$, the density operator can be expressed in terms of the bright modes as
$\begin{split}&\hat{n}_{\bm{q}_{\parallel}}=\sum_{n^{\prime},\lambda^{\prime}\in\mathcal{N}_{\rm unocc}}\sum_{n,\lambda\in\mathcal{N}_{\rm occ}}e^{-q_{\parallel}^{2}\ell_{B}^{2}/4}\sqrt{N_{\rm L}}\times\\ &\times\Big{[}\mathcal{F}_{n^{\prime},\lambda^{\prime},n,\lambda}(\bm{q}_{\parallel})\hat{b}^{\dagger}_{n,\lambda,n^{\prime},\lambda^{\prime}}(q_{x})+\\ &+\mathcal{F}_{n,\lambda,n^{\prime},\lambda^{\prime}}(\bm{q}_{\parallel})\hat{b}_{n,\lambda,n^{\prime},\lambda^{\prime}}(-q_{x})\Big{]}~{}.\end{split}$ (79)
In this form, the density operator is expressed as a sum of creation and annihilation operators for the bright modes. By substituting the expression from Eq. (LABEL:eq:bright_density) into the interaction Hamiltonian and using the property $\phi_{-\bm{q}_{\parallel},n_{z},s}(0)=\phi_{\bm{q}_{\parallel},n_{z},s}(0)$, we obtain:
^$\begin{split}\hat{\mathcal{H}}_{\rm int}&=-\sum_{\bm{q}_{\parallel},n_{z},\alpha,\beta}\sum_{n^{\prime},\lambda^{\prime}\in\mathcal{N}_{\rm unocc}}\sum_{n,\lambda\in\mathcal{N}_{\rm occ}}\Big{(}\hat{a}^{\dagger}_{\bm{q}_{\parallel},n_{z}}+\hat{a}_{-\bm{q}_{\parallel},n_{z}}\Big{)}\times\\ &\times\Big{[}\hbar g_{\bm{q}_{\parallel},n_{z},s,n,\lambda,n^{\prime},\lambda^{\prime}}\hat{b}_{n,\lambda,n^{\prime},\lambda^{\prime}}(q_{x})+\\ &+\hbar g_{\bm{q}_{\parallel},n_{z},s,n^{\prime},\lambda^{\prime},n,\lambda}\hat{b}_{n,\lambda,n^{\prime},\lambda^{\prime}}(-q_{x})\Big{]}~{},\end{split}$ (80)
where the couplings are defined as
$g_{\bm{q}_{\parallel},n_{z},s,n,\lambda,n^{\prime},\lambda^{\prime}}\equiv\frac{e}{\hbar}\phi_{{q}_{\parallel},n_{z},s}(0)\sqrt{\frac{N_{\rm L}}{S}}e^{-q_{\parallel}^{2}\ell_{B}^{2}/4}\mathcal{F}_{n,\lambda,n^{\prime},\lambda^{\prime}}(\bm{q}_{\parallel})~{},$ (81)
and the Rabi frequencies are defined as the modulus of the coupling terms,
$\Omega_{\bm{q}_{\parallel},n_{z},s,n,\lambda,n^{\prime},\lambda^{\prime}}=|g_{\bm{q}_{\parallel},n_{z},s,n,\lambda,n^{\prime},\lambda^{\prime}}|~{}.$ (82)
Notice that the interaction Hamiltonian in Eq. (80) is quadratic in the bosonic operators $\hat{a}^{\dagger}_{\bm{q}_{\parallel},n_{z}},\hat{a}_{-\bm{q}_{\parallel},n_{z}},\hat{b}_{n,\lambda,n^{\prime},\lambda^{\prime}}(-q_{x}),\hat{b}_{n,\lambda,n^{\prime},\lambda^{\prime}}(q_{x})$.
We now focus on the Coulomb term
^$\hat{\mathcal{H}}_{\rm C}=\frac{1}{2S}\sum_{\bm{q}_{\parallel}}\hat{n}_{-\bm{q}_{\parallel}}e^{2}g_{\infty}^{(0)}(q_{\parallel},0,0)\hat{n}_{\bm{q}_{\parallel}}~{},$ (83)
where $g_{\infty}^{(0)}(\bm{q}_{\parallel},0,0)$ is given in Eq. (LABEL:eqn:green_function_hBN_classical_inf). As the density operator $\hat{n}_{\bm{q}_{\parallel}}$, as expressed in Eq. (LABEL:eq:bright_density), is linear in the bright modes, the Coulomb term is thus quadratic in the bright modes.
At this point, all the Hamiltonian terms in Eq. (66) are quadratic in terms of the bright and cavity modes $\hat{b}^{\dagger}_{n,\lambda,n^{\prime},\lambda^{\prime}}(q_{x}),\hat{b}_{n,\lambda,n^{\prime},\lambda^{\prime}}(q_{x}),\hat{a}^{\dagger}_{\bm{q}_{\parallel},n_{z},s},\hat{a}_{\bm{q}_{\parallel},n_{z},s}$. Hence, the Hamiltonian $\hat{\mathcal{H}}$ in Eq. (66) can be diagonalized by a Bogoliubov transformation. The resulting eigenfrequencies, denoted as ${\omega}_{\nu}$, describe the energies of the excitation modes in the system. Here, the index $\nu$ encapsulates all relevant quantum numbers. Such eigenfrequencies are given by
$\begin{split}1-e^{2}g^{(0)}({q}_{\parallel},0,0,\omega_{\nu})\chi^{(0)}_{\rm nn}({q}_{\parallel},\omega_{\nu})=0~{},\end{split}$ (84)
corresponding to setting to zero the RPA dielectric function in Eq. (5), $\epsilon_{\rm RPA}({q}_{\parallel},\omega_{\nu})=0$.
V.2 The vacuum Rabi frequencies
[…]
Figure 7: (Color online) Panels (a,d) display the ratio $R_{q_{\parallel},n_{z},s,-,0,+,1}=\Omega_{q_{\parallel},n_{z},s,-,0,+,1}/\omega_{q_{\parallel},n_{z},s}$, where $\Omega_{q_{\parallel},n_{z},s,-,0,+,1}$ denotes the Rabi frequency for the Landau level (LL) transition $(0,+1)$, and $\omega_{q_{\parallel},n_{z},s}$ is the cavity mode frequency. This ratio is shown as a function of $q_{\parallel}$ (in ${\rm nm}^{-1}$) for different values of $n_{z}$, with $L_{z}=30{\rm nm}$ and $B=10{\rm T}$. Panels (b,e) display the same ratio as a function of $L_{z}$ (in ${\rm nm}$) for different $n_{z}$ for a fixed $\bar{q}_{\parallel}=0.1{\rm nm^{-1}}$ and $B=10{\rm T}$. Panels (c,f) present the Rabi frequency as a function of $\hbar v_{\rm F}/\ell_{B}$ (in ${\rm eV}$) for different $n_{z}$. Panels (a,b,c) show these quantities for the upper Reststrahlen band using blue dotted curves, while panels (d,e,f) present the lower Reststrahlen band with red dashed curves. The line thickness diminishes as the quantum number $n_{z}$ increases.
We now focus on the real part of the dielectric function. The real part of the scalar Green function in Eq. (LABEL:eqn:green_function_hBN_expansion_quantum), for $z=z^{\prime}=0$, reads
${\rm Re}~{}g_{\rm q}^{(0)}({q}_{\parallel},0,0,\omega)=\sum_{n_{z},s}\phi^{2}_{{q}_{\parallel},n_{z},s}({0})\left(\frac{2\hbar\omega_{q_{\parallel},n_{z},s}}{(\hbar\omega)^{2}-(\hbar\omega_{q_{\parallel},n_{z},s})^{2}}\right)~{}.$ (85)
[…]
Figure 8: (Color online) Panel (a) shows $-{\rm Im}~{}{\chi}^{\rm st}_{\rm nn}({q}_{\parallel},\omega)$ as a function of the wavevector $q_{\parallel}$ for a given magnetic field, $B=5~{}{\rm T}$, and for $L_{z}=50~{}{\rm nm}$. Panel (b) shows the same quantity in the absence of the cavity confinement $L_{z}\to\infty$. Panel (c) shows $-{\rm Im}~{}{\chi}_{\rm nn}({q}_{\parallel},\omega)$ as a function of $\hbar v_{\rm F}/\ell_{B}$, for fixed momentum $\bar{q}_{\parallel}=0.05~{}{\rm nm}^{-1}$ and for $L_{z}=50~{}{\rm nm}$. Other parameters are $N_{\rm c}=30$, $n_{z}^{\rm max}=30$, $\hbar\eta=5\times 10^{-4}~{}{\rm meV}$.
The real part of the density–density response function, given in Eq. (65) is
$\begin{split}{\rm Re}\chi^{(0)}_{nn}(q_{\parallel},\omega)=\sum_{n,\lambda\in\mathcal{N}_{\rm occ}}\sum_{n^{\prime},\lambda^{\prime}\in\mathcal{N}_{\rm unocc}}\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})\times\\ \times\left(\frac{2(\varepsilon_{\lambda^{\prime},n^{\prime}}-\varepsilon_{\lambda,n})}{(\hbar\omega)^{2}-(\varepsilon_{\lambda,n}-\varepsilon_{\lambda^{\prime},n^{\prime}})^{2}}\right)~{}.\end{split}$ (86)
Thus, the real part of the dielectric function can be conveniently written as
${\rm Re}~{}\epsilon_{\rm RPA}(q_{\parallel},\omega)=1-\sum_{n_{z},s}\sum_{n,\lambda\in\mathcal{N}_{\rm occ}}\sum_{n^{\prime},\lambda^{\prime}\in\mathcal{N}_{\rm unocc}}{(\hbar\Omega_{q_{\parallel},{n}_{z},{s},{\lambda},{n},{\lambda}^{\prime},{n}^{\prime}})^{2}}\left(\frac{2\hbar\omega_{q_{\parallel},n_{z},s}}{\left[(\hbar\omega)^{2}-(\hbar\omega_{q_{\parallel},n_{z},s})^{2}\right]}\frac{{2(\varepsilon_{\lambda^{\prime},n^{\prime}}-\varepsilon_{\lambda,n})}}{\left[(\hbar\omega)^{2}-(\varepsilon_{\lambda,n}-\varepsilon_{\lambda^{\prime},n^{\prime}})^{2}\right]}\right)~{},$ (87)
where the Rabi frequencies $\Omega_{q_{\parallel},{n}_{z},{s},{\lambda},{n},{\lambda}^{\prime},{n}^{\prime}}$ are defined as
$\hbar\Omega_{q_{\parallel},{n}_{z},{s},{\lambda},{n},{\lambda}^{\prime},{n}^{\prime}}=\phi_{{q}_{\parallel},n_{z},s}(0)\sqrt{{\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})}}~{}.$ (88)
Eq. (87) can be interpreted as describing the response of a system composed of harmonic modes with frequencies $\omega_{q_{\parallel},n_{z},s}$, which are coupled to other harmonic modes with frequencies $(\varepsilon_{\lambda^{\prime},n^{\prime}}-\varepsilon_{\lambda,n})/\hbar$. The coupling strength between these modes is characterized by the Rabi frequency $\hbar\Omega_{q_{\parallel},{n}z,{s},{\lambda},{n},{\lambda}^{\prime},{n}^{\prime}}$.
In standard C-QED dalibard1982 ; DDB_arxiv_2023 ; ciuti2005 , it is well known that the Rabi frequency is proportional to the product of the vacuum fluctuations of the electric field $\sqrt{\langle\hat{E}^{2}\rangle}$ and the dipole moment of the atom in the cavity, $d$, such that $\Omega_{\rm R}\sim d\sqrt{\langle\hat{E}^{2}\rangle}$. Eq. (88) generalizes this concept to non-local ($q_{\parallel}\neq 0$), sub-wavelength cavities interacting with a 2D material. As seen from Eq. (88), the Rabi frequency is proportional to the product of the vacuum fluctuation coefficients $\phi_{{q}_{\parallel},n_{z},s}(0)$ and the density matrix elements $\sqrt{{\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})}}$. Therefore, to maximize the light-matter coupling, one can separately optimize the vacuum fluctuations and the density matrix elements. Specifically, $\phi_{{q}_{\parallel},n_{z},s}(0)$ reaches its maximum when $q_{\parallel}L_{z}\sim 1$, while $\sqrt{{\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})}}$ peaks when $q_{\parallel}\ell_{B}\sim 1$. This leads to the conclusion that for optimal coupling, all the relevant length scales must be comparable, such that $q_{\parallel}\sim 1/\ell_{B}\sim 1/L_{z}$.
In Fig. 7, we present the ratio $R_{q_{\parallel},n_{z},s,-,0,+,1}\equiv(\Omega_{q_{\parallel},n_{z},s,-,0,+,1}/\omega_{q_{\parallel},n_{z},s})$, where $\Omega_{q_{\parallel},n_{z},s,-,0,+,1}$ is the Rabi frequency associated with the LL transition $(0,+1)$ and $\omega_{q_{\parallel},n_{z},s}$ is the cavity frequency as a function of different parameters. In Fig. 7(a,d), the ratio $R_{q_{\parallel},n_{z},s,-,0,+,1}$ is plotted as a function of $q_{\parallel}$ (in ${\rm nm}^{-1}$) for various $n_{z}$. This plot reveals a non-monotonic behavior in $q_{\parallel}$, where each mode exhibits a specific $q_{\parallel}$ that maximizes the Rabi frequency. Similarly, Fig. 7(b,e) displays this quantity as a function of $L_{z}$ (in ${\rm nm}$) for different $n_{z}$. This also shows a non-monotonic dependence on $L_{z}$, which contrasts with the naive expectation that the Rabi frequency would scale as $1/\sqrt{L_{z}}$. This scaling is valid only for $q_{\parallel}L_{z}\gg 1$, whereas for small $L_{z}$, the coupling vanishes. Finally, Fig. 7(c,f) shows the Rabi frequency as a function of $\hbar v_{\rm F}/\ell_{B}$, which can be controlled experimentally by adjusting the external magnetic field. At low values, specifically around $\hbar v_{\rm F}/\ell_{B}\approx 0.01~{}{\rm eV}$, increasing the magnetic field—and thus this ratio—leads to a stronger coupling. However, the coupling saturates when $\hbar v_{\rm F}/\ell_{B}$ reaches approximately $0.15~{}{\rm eV}$ or higher.
In Fig. 7(a-f), we observe that the maximum ratio $R_{q_{\parallel},n_{z},s,-,0,+,1}$, achieved for $n_{z}=0$, exceeds $10\%$ for the upper Reststrahlen band and approaches $10\%$ for the lower band. This ratio places the system firmly within the ultra-strong coupling regime, where the coupling strength is a significant fraction of the cavity frequency Kockum19 . Additionally, a single LL hybridizes with multiple cavity modes, as the Rabi frequencies of many different modes remain non-negligible. This situation characterizes a regime known as super-strong coupling kuzmin2019 .
V.3 The depolarization shift
Fig. 3(a) shows $-{\rm Im}~{}{\chi}_{\rm nn}({q}_{\parallel},\omega)$ as a function of the wavevector $q_{\parallel}$ for a fixed magnetic field, $B=5~{}{\rm T}$. This plot reveals that, outside the Reststrahlen bands, LL excitations exhibit a $q_{\parallel}$-dependent positive frequency shift, known as the depolarization shift Giuliani_and_Vignale ; shtrichman2001 . Here, we investigate the origin of this effect. First, we recall that the dressed density–density response function is defined as ${\chi}_{\rm nn}({q}_{\parallel},\omega)={\chi}^{(0)}_{\rm nn}({q}_{\parallel},\omega)/\epsilon_{\rm RPA}({q}_{\parallel},\omega)$. We further define the static dressed density–density response function ${\chi}^{\rm st}_{\rm nn}({q}_{\parallel},\omega)$ as the response function dressed by $\epsilon^{\rm st}_{\rm RPA}({q}_{\parallel},\omega)$, which considers only the static part of the Green function $g_{\infty}^{(0)}(\bm{r},\bm{r}^{\prime})$, excluding contributions from quantum modes:
$\begin{split}{\chi}^{\rm st}_{\rm nn}({q}_{\parallel},\omega)&=\frac{{\chi}^{(0)}_{\rm nn}({q}_{\parallel},\omega)}{\epsilon^{\rm st}_{\rm RPA}({q}_{\parallel},\omega)}~{},\\ {\epsilon^{\rm st}_{\rm RPA}({q}_{\parallel},\omega)}&=1-e^{2}g_{\infty}^{(0)}({q}_{\parallel},0,0)\chi^{(0)}_{\rm nn}({q}_{\parallel},\omega)~{}.\\ \end{split}$ (89)
Fig. 8(a) displays $-{\rm Im}~{}{\chi}^{\rm st}_{\rm nn}({q}_{\parallel},\omega)$ as a function of the wavevector $q_{\parallel}$ for $B=5~{}{\rm T}$. Comparing this plot with Fig. 3(a), we observe that the off-resonant features outside the Reststrahlen bands, such as the diamagnetic shift, are accurately captured by the static approximation $-{\rm Im}~{}{\chi}^{\rm st}_{\rm nn}({q}_{\parallel},\omega)$. However, the pronounced dispersion of the LL transitions is a cavity-induced effect. This becomes evident in Fig. 8(b), where we plot $-{\rm Im}~{}{\chi}^{\rm st}_{\rm nn}({q}_{\parallel},\omega)$ in the limit of no cavity confinement ($L_{z}\to\infty$). In this limit, the depolarization shifts are significantly reduced, and the LL transitions become almost flat. This indicates that the diamagnetic shift arises from the Coulomb propagator $g_{\infty}^{(0)}({q}_{\parallel},0,0)$, which is strongly influenced by the cavity confinement $L_{z}$, as described by Eq. (LABEL:eqn:green_function_hBN_classical_inf).
V.4 Two-transition model
Fig. 8(c) shows $-{\rm Im}~{}{\chi}_{\rm nn}(q_{\parallel},\omega)$ as a function of $\hbar v_{\rm F}/\ell_{B}$, for a small wavevector $\bar{q}_{\parallel}=0.05~{}{\rm nm}^{-1}$ and $L_{z}=50~{}{\rm nm}$. In this regime, where $q_{\parallel}L_{z}\gg 1$, only the Rabi frequency with $n_{z}=0$ remains significant, as the contributions from other $n_{z}$ modes rapidly decay to zero, as evident in Fig. 7(a). This behavior allows us to simplify the analysis by focusing exclusively on the interaction between a specific cavity mode $(q_{\parallel},\bar{n}_{z},\bar{s})$ and a single LL transition$(\bar{n},\bar{\lambda})\rightarrow(\bar{n}^{\prime},\bar{\lambda}^{\prime})$. Under this approximation, the dielectric function ${\rm Re}~{}\epsilon^{(2)}_{\rm RPA}(q_{\parallel},\omega)$ can be expressed as:
$\begin{split}&{\rm Re}~{}\epsilon^{(2)}_{\rm RPA}(q_{\parallel},\omega)\equiv\\ &=1-\frac{4(\hbar\Omega_{q_{\parallel},\bar{n}_{z},\bar{s},\bar{\lambda},\bar{n},\bar{\lambda}^{\prime},\bar{n}^{\prime}})^{2}{(\varepsilon_{\bar{\lambda}^{\prime},\bar{n}^{\prime}}-\varepsilon_{\bar{\lambda},\bar{n}})}\hbar\omega_{q_{\parallel},\bar{n}_{z},\bar{s}}}{\left[(\hbar\omega)^{2}-(\hbar\omega_{q_{\parallel},\bar{n}_{z},\bar{s}})^{2}\right]\left[(\hbar\omega)^{2}-(\varepsilon_{\bar{\lambda},\bar{n}}-\varepsilon_{\bar{\lambda}^{\prime},\bar{n}^{\prime}})^{2}\right]}~{}.\end{split}$ (90)
where all quantum numbers $x$ are fixed and denoted by $\bar{x}$. The corresponding frequencies of the coupled modes can be derived by solving the quadratic equation, obtained by setting ${\rm Re}~{}\epsilon^{(2)}_{\rm RPA}(q_{\parallel},\omega)=0$:
$\begin{split}&(\hbar\omega)^{2}=\frac{(\hbar\omega_{q_{\parallel},\bar{n}_{z},\bar{s}})^{2}+(\varepsilon_{\bar{\lambda},\bar{n}}-\varepsilon_{\bar{\lambda}^{\prime},\bar{n}^{\prime}})^{2}}{{2}}+\\ &\pm\frac{\sqrt{\left[(\hbar\omega_{q_{\parallel},\bar{n}_{z},\bar{s}})^{2}-(\varepsilon_{\bar{\lambda},\bar{n}}-\varepsilon_{\bar{\lambda}^{\prime},\bar{n}^{\prime}})^{2}\right]^{2}+16(\hbar\Omega_{q_{\parallel},\bar{n}_{z},\bar{s},\bar{\lambda},\bar{n},\bar{\lambda}^{\prime},\bar{n}^{\prime}})^{2}\hbar\omega_{q_{\parallel},\bar{n}_{z},\bar{s}}(\varepsilon_{\bar{\lambda},\bar{n}}-\varepsilon_{\bar{\lambda}^{\prime},\bar{n}^{\prime}})}}{2}~{}.\end{split}$ (91)
These frequencies are plotted as orange dotted lines in Fig. 8(c) for the $(0,1)$ LL transition coupled to the hPP mode with $n_{z}=0$ and $s=+$. The bare LL transitions are shown as white lines. As evident from the figure, this simplified model fails to capture the correct excitation spectrum of the system accurately. Even when the two bare excitations are detuned, the model only correctly predicts the cavity mode, while it fails to reproduce the shifted LL transition, which experiences a positive depolarization shift. This shift arises due to the influence of the static Coulomb potential $g_{\infty}^{(0)}({q}_{\parallel},0,0)$.
Therefore, to achieve a more accurate description of the spectrum, it is essential to include the static Coulomb potential $g_{\infty}^{(0)}({q}_{\parallel},0,0)$ in the analysis, which modifies Eq. (90) as follows
$\begin{split}&{\rm Re}~{}\epsilon^{(2+{\rm C})}_{\rm RPA}(q_{\parallel},\omega)=\\ &=1-\frac{4(\hbar\Omega_{q_{\parallel},\bar{n}_{z},\bar{s},\bar{\lambda},\bar{n},\bar{\lambda}^{\prime},\bar{n}^{\prime}})^{2}{(\varepsilon_{\bar{\lambda}^{\prime},\bar{n}^{\prime}}-\varepsilon_{\bar{\lambda},\bar{n}})}\hbar\omega_{q_{\parallel},\bar{n}_{z},\bar{s}}}{\left[(\hbar\omega)^{2}-(\hbar\omega_{q_{\parallel},\bar{n}_{z},\bar{s}})^{2}\right]\left[(\hbar\omega)^{2}-(\varepsilon_{\bar{\lambda},\bar{n}}-\varepsilon_{\bar{\lambda}^{\prime},\bar{n}^{\prime}})^{2}\right]}+\\ &-e^{2}g_{\infty}^{(0)}({q}_{\parallel},0,0)\left(\frac{2(\varepsilon_{\bar{\lambda}^{\prime},\bar{n}^{\prime}}-\varepsilon_{\bar{\lambda},\bar{n}})\mathcal{M}_{\bar{n}^{\prime},\bar{\lambda}^{\prime},\bar{n},\bar{\lambda}}({q}_{\parallel})}{(\hbar\omega)^{2}-(\varepsilon_{\bar{\lambda},\bar{n}}-\varepsilon_{\bar{\lambda}^{\prime},\bar{n}^{\prime}})^{2}}\right)~{}.\end{split}$
The solutions of ${\rm Re}~{}\epsilon^{(2+{\rm C})}_{\rm RPA}(q_{\parallel},\omega)$ are shown as red solid lines in Fig. 8(c) for the $(0,1)$ LL transition coupled to the hPP mode with $n_{z}=0$ and $s=+1$. These solutions accurately reproduce the observed spectrum, demonstrating both the critical role of the coupling to the resonant mode with $n_{z}=0$ and $s=+$ and the importance of incorporating the static Coulomb potential $g_{\infty}^{(0)}({q}_{\parallel},0,0)$ in the model. In this case, the Rabi frequency is approximately $\Omega_{q_{\parallel},0,z,+1,-,0,+,1}\approx 0.014~{}{\rm eV}$, which corresponds to about $7\%$ of the bare cavity frequency. This value indicates that the system, for low wavevectors, is near the threshold of the ultra-strong coupling regime.
VI Conclusions
Here, we investigated the polaritonic spectrum of LLs in graphene embedded in an HPP cavity footnote , including non-local effects. Our analysis demonstrated that the non-local response introduces spectral features absent in purely local theories, such as coupling to LL transitions that are dipole-forbidden, underscoring its crucial role in accurately describing light-matter interactions at the nanoscale. Notably, our nonlocal theory identifies the vacuum Rabi frequency is maximized when the light’s momentum, $q_{\parallel}$, is comparable to the inverse of the magnetic length, $q_{\parallel}\ell_{B}\sim 1$. This insight underscores the importance of considering nonlocal effects for achieving optimal light-matter coupling in such systems. Finally, we find that the strong coupling between LL transitions in graphene and cavity modes can be effectively engineered under high magnetic fields and low temperatures. Ref. andolina2024, established that sub-wavelength cavities are described by a scalar potential $\phi$ rather then a vector potential. Starting from this insight, we developed a new theoretical framework to study C-QED of sub-wavelength cavities, by quantizing the scalar potential $\hat{\phi}$, assessing quantum vacuum fluctuations, and providing a way to calculate vacuum Rabi frequencies, without resorting to phenomenological compression factors. Our theory provides a quantitative tool to directly assess light-matter interactions in the strong and ultra-strong coupling regimes, offering a predictive tool for cavity design.
Recent studies dapolito2023 ; wehmeier2024 have investigated the coupling between hBN phonon-polaritons and LL transitions in graphene in a SNOM setup, showing evidence of strong-coupling. On the other hand, polaritonic engineering HerzigSheinfux stands as a promising tool for designing light-matter interaction at the nanoscale in van der Waals materials, and studying the strong-coupling regime at sub-Kelvin temperature, in the presence of strong magnetic fields.
Our approach would open new avenues for engineering the electronic properties of materials embedded in cavities. For example, it could enable the design of superconductivity in moiré materials through tailored cavity structures BlochReview , the modulation of topological phenomena such as the quantum Hall effect appugliese2022 ; enkner2024 , or the enhancement and control of interaction effects, including fractional quantum Hall states enkner2024b ; bacciconi2023 or correlated states in twisted bilayer graphene Cao in a cavity. Indeed, our microscopic theory of sub-wavelength cavities can be used to investigate how cavities can be employed to enhance Ebbesen19 ; Sentef_ScienceAdvances_2018 or induce schlawin2019 superconductivity.
Another promising direction for future research would be the study of polaritonic lattices Jacqmin2014 , wherein HPP cavities are constructed from a lattice of different materials or incorporate a grating grown on the cavity, thereby breaking strict translational symmetry. In such systems, energy bands naturally emerge and can be described using Bloch theory. This additional degree of control could significantly enhance the tunability of polaritonic systems. Our theoretical framework could be readily generalized to describe these more complex lattice-based systems.
VI.1 Acknowledgments
G.M.A acknowledges useful discussion with I. Torre, L. Orsini, J. Sueiro, G. Mazza, Z. Bacciconi, D. De Bernardis, G. Chiriacò, M. Michael, A. Rubio, M. Ruggenthaler and F. Schlawin. G.M.A acknowledge funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie (Grant Agreement No. 101146870 – COMPASS). G.M.A and M.S. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101002955 – CONQUER). B.T. acknowledges funding from “la Caixa” Foundation (ID 100010434). R.R. and M.P. were supported by the MUR - Italian Ministry of University and Research under the “Research projects of relevant national interest - PRIN 2020” - Project No. 2020JLZ52N (“Light-matter interactions and the collective behavior of quantum 2D materials, q-LIMA”) and by the European Union under grant agreement No. 101131579 - Exqiral and No. 873028 - Hydrotronics. F.H.L.K. acknowledges support from the ERC TOPONANOP under grant agreement no. 726001, the Gordon and Betty Moore Foundation through Grant GBMF12212, and the Government of Spain (FIS2016-81044; PID2019-106875GB-100; Severo Ochoa CEX2019-000910-S [MCIN/ AEI/10.13039/501100011033], PCI2021-122020-2A, and PDC2022-133844-100 funded by MCIN/AEI/10.13039/501100011033). This work was also supported by the European Union NextGenerationEU/PRTR (PRTR-C17.I1) and EXQIRAL 101131579, Fundació Cellex, Fundació Mir-Puig, and Generalitat de Catalunya (CERCA, AGAUR, 2021 SGR 014431656). Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Additionally, the research leading to these results received funding from the European Union’s Horizon 2020 under grant agreements no. 881603 (Graphene flagship Core3) and 820378 (Quantum flagship). This material is based upon work supported by the Air Force Office of Scientific Research under award number FA8655-23-1-7047. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the United States Air Force.
Appendix A The dielectric tensor in hBN
hBN is a two-dimensional material composed of boron and nitrogen atoms arranged in a hexagonal lattice structure, akin to graphene. Here, we consider stacked layers of hBN, forming a three-dimensional structure with weak van der Waals interactions between the layers. This material exhibits hyperbolic behavior due to its anisotropic dielectric tensor $\bm{\epsilon}(\bm{r},\omega)={\rm diag}(\epsilon_{x}(\bm{r},\omega),\epsilon_{x}(\bm{r},\omega),\epsilon_{z}(\bm{r},\omega))$, where the real parts of its principal components have opposite signs in the Reststrahlen bands frequency ranges. The components of the dielectric tensor are given by the Drude-Lorentz formula tomadin2015 :
$\epsilon_{i}(\omega)=\epsilon_{i,\infty}+\frac{\epsilon_{i,0}-\epsilon_{i,\infty}}{1-(\omega/\omega^{\rm T}_{i})^{2}}~{},$ (92)
where $i=x,z$ and $\omega^{\rm T}_{i}$ are the transverse phonon frequencies, $\epsilon_{i,0}$ is the static dielectric constant ($\epsilon_{i}(0)=\epsilon_{i,0}$) while $\epsilon_{i,\infty}$ is the high-frequency limit, $\epsilon_{i}(\omega\to\infty)=\epsilon_{i,\infty}$. The longitudinal phonon frequencies, defined by $\epsilon_{i}(\omega)$ can be obtained by the relation $\omega^{\rm L}_{i}=\omega^{\rm T}_{i}\sqrt{(\epsilon_{i,0}/\epsilon_{i,\infty})}$. Realistic values for hBN of these quantities Geick66 are reported in Tab. 3.
$i=x$ $i=z$
$\epsilon_{i,\infty}$ 4.5 2.5
$\epsilon_{i,0}$ 6.4 3.0
$\omega^{\rm T}_{i}$ (meV) 168 94
$\omega^{\rm L}_{i}$ (meV) 200 103
Table 3: Realistic values of the parameters for the hBN dielectric tensor Geick66 entering in the Drude-Lorentz formula in Eq. (92).
Small dissipation can be included in the dielectric tensor in Eq. (93) by performing the substitution $\omega\to\omega+\mathrm{i}\eta$ and getting:
$\epsilon_{i}(\omega)\approx\epsilon_{i,\infty}+\frac{\epsilon_{i,0}-\epsilon_{i,\infty}}{1-(\omega/\omega^{\rm T}_{i})^{2}-\mathrm{i}2\omega\eta/(\omega^{\rm T}_{i})^{2}}~{},$ (93)
valid for $\eta\to 0$.
[…]
Figure 9: (Color online) The in-plane (out-of-plane) dielectric function $\epsilon_{x}(\omega)$ ($\epsilon_{z}(\omega)$) is plotted as a function of frequency $\omega$, using the realistic values reported in Table 3. The two Reststrahlen bands, where $\epsilon_{x}(\omega)\epsilon_{z}(\omega)<0$, are highlighted in gray.
Appendix B The density–density response function for graphene in a magnetic field
Here, we calculate the bare density-density response function, $\chi^{(0)}_{nn}(q_{\parallel},\omega)$. First of all, it is useful to define the ladder operators $\hat{a}$ and $\hat{a}^{\dagger}$ for the case when the displacement $k_{x}\ell^{2}_{B}$ is zero, as:
$\begin{split}\hat{{y}}&=\frac{\ell_{B}}{\sqrt{2}}\left(\hat{a}+\hat{a}^{\dagger}\right)~{},\\ \hat{p}_{y}&=\frac{\hbar\mathrm{i}}{\sqrt{2}\ell_{B}}\left(\hat{a}^{\dagger}-\hat{a}\right)~{}.\end{split}$ (94)
It is now useful to introduce the displacement operators defined as:
$\begin{split}\hat{\mathcal{D}}\left(\alpha\right)&\equiv\exp\left(\alpha\hat{a}^{\dagger}-\alpha^{*}\hat{a}\right)~{}\end{split}$ (95)
where $\alpha$ is a complex number. The displaced ladder operators $\hat{a}_{k_{x}}$ and $\hat{a}_{k_{x}}^{\dagger}$ can be related to the ladder operators $\hat{a}$ and $\hat{a}^{\dagger}$ by introducing the displacement operator $\hat{\mathcal{D}}\left(\frac{k_{x}\ell_{B}}{\sqrt{2}}\right)$ as:
$\begin{split}\hat{\mathcal{D}}\left(\frac{k_{x}\ell_{B}}{\sqrt{2}}\right)&\equiv\exp\left(\frac{k_{x}\ell_{B}}{\sqrt{2}}\left(\hat{a}^{\dagger}-\hat{a}\right)\right)~{},\\ &=\exp\left(-\mathrm{i}k_{x}\ell_{B}^{2}\hat{p}_{y}/\hbar\right)~{}.\end{split}$ (96)
The operator $\hat{\mathcal{D}}\left(\frac{k_{x}\ell_{B}}{\sqrt{2}}\right)$ links the ladder operators $\hat{a}_{k_{x}}$ with $\hat{a}$:
$\begin{split}\hat{a}_{k_{x}}&=\hat{\mathcal{D}}\left(\frac{k_{x}\ell_{B}}{\sqrt{2}}\right)~{}\hat{{a}}~{}\hat{\mathcal{D}}^{\dagger}\left(\frac{k_{x}\ell_{B}}{\sqrt{2}}\right)~{},\\ &=\hat{a}-\frac{k_{x}\ell_{B}}{\sqrt{2}}~{}.\end{split}$ (97)
Hence, the operator $\hat{\mathcal{D}}\left(\frac{k_{x}\ell_{B}}{\sqrt{2}}\right)$ allows us to connect $\hat{y}_{k_{x}}$ and $\hat{y}$:
$\begin{split}\hat{{y}}_{k_{x}}&=\hat{\mathcal{D}}\left(\frac{k_{x}\ell_{B}}{\sqrt{2}}\right)~{}\hat{{y}}~{}\hat{\mathcal{D}}^{\dagger}\left(\frac{k_{x}\ell_{B}}{\sqrt{2}}\right)~{},\\ &=\hat{{y}}-{k_{x}}{\ell_{B}^{2}}~{}.\end{split}$ (98)
Hence, by using the displacement operator $\hat{\mathcal{D}}\left(\frac{k_{x}\ell_{B}}{\sqrt{2}}\right)$, we can relate the displaced vacuum |$\ket{0_{k_{x}}}$ (defined by $\hat{a}_{k_{x}}\ket{0_{k_{x}}}=0$) with the vacuum |$\ket{0}$, such that $\hat{a}\ket{0}=0$:
|$\ket{0_{k_{x}}}$ $=\hat{\mathcal{D}}\left(\frac{k_{x}\ell_{B}}{\sqrt{2}}\right)\ket{0}~{}.$ (99)
Similarly, we can link the excited states:
|$\ket{n_{k_{x}}}$ $=\hat{\mathcal{D}}\left(\frac{k_{x}\ell_{B}}{\sqrt{2}}\right)\ket{n},$ (100)
where |$\ket{n}$ is the Fock state with respect to $\hat{a}$, |$\ket{n}=\left[{(\hat{a}^{\dagger})^{n}}/{\sqrt{n!}}\right]\ket{0}$
We now alculate the matrix elements of the density operator $\hat{n}_{\bm{q}_{\parallel}}\equiv\exp(-\mathrm{i}\bm{q}_{\parallel}\cdot\bm{r}_{\parallel})$.
$\braket{{k_{x}}^{\prime}}{e^{-\mathrm{i}q_{x}\hat{x}}}{{k_{x}}}=\delta_{k_{x}^{\prime},k_{x}+q_{x}}~{},$ (101)
$\begin{split}&\braket{n^{\prime}_{{k_{x}}^{\prime}}}{e^{-\mathrm{i}q_{y}\hat{y}}}{n_{k_{x}}}=\\ &=\braket{n^{\prime}}{\hat{\mathcal{D}}^{\dagger}\left(\frac{k_{x}^{\prime}\ell_{B}}{\sqrt{2}}\right)\hat{\mathcal{D}}^{\dagger}\left(\frac{\mathrm{i}q_{y}\ell_{B}}{\sqrt{2}}\right)\hat{\mathcal{D}}\left(\frac{k_{x}\ell_{B}}{\sqrt{2}}\right)}{n}~{},\end{split}$ (102)
where we used that $e^{-\mathrm{i}q_{y}\hat{y}}=\hat{\mathcal{D}}^{\dagger}(\mathrm{i}q_{y}\ell_{B}/\sqrt{2})$ and applied Eq. (100) to connect the states |$\ket{n^{\prime}_{k_{x}^{\prime}}}$ with |$\ket{n^{\prime}}$ and |$\ket{n_{k_{x}}}$ with |$\ket{n}$. Using the braiding relation $\hat{\mathcal{D}}^{\dagger}(\alpha)\hat{\mathcal{D}}^{\dagger}(\beta)=\exp(\alpha^{*}\beta-\alpha\beta^{*})\hat{\mathcal{D}}^{\dagger}(\beta)\hat{\mathcal{D}}^{\dagger}(\alpha)$ we get:
$\begin{split}&\braket{n^{\prime}_{{k_{x}}^{\prime}}}{e^{-\mathrm{i}q_{y}\hat{y}}}{n_{k_{x}}}=\\ &=\exp(-\mathrm{i}k^{\prime}_{x}q_{y}\ell_{B}^{2})\braket{n^{\prime}}{\hat{\mathcal{D}}^{\dagger}\left(\frac{[(k_{x}^{\prime}-k_{x})+\mathrm{i}q_{y}]\ell_{B}}{\sqrt{2}}\right)}{n}~{},\end{split}$ (103)
Employing Eqs. (101,103) we can express the element matrix of the density operator $\hat{n}_{\bm{q}_{\parallel}}$ as:
$\begin{split}&\braket{{k_{x}}^{\prime},n^{\prime}}{\hat{n}_{\bm{q}_{\parallel}}}{{k_{x}},n}=\\ &=\exp(-\mathrm{i}k^{\prime}_{x}q_{y}\ell_{B}^{2})\delta_{k_{x}^{\prime},k_{x}+q_{x}}\braket{n^{\prime}}{\hat{\mathcal{D}}^{\dagger}\left(\frac{q_{z}\ell_{B}}{\sqrt{2}}\right)}{n}~{}.\end{split}$ (104)
where we defined $q_{z}\equiv q_{x}+\mathrm{i}q_{y}$. By using the Baker–Campbell–Hausdorff formula for $[\hat{A},[\hat{A},\hat{B}]]=[\hat{B},[\hat{A},\hat{B}]]=0$,
$e^{\hat{A}+\hat{B}}=e^{\hat{A}}e^{\hat{B}}e^{-[\hat{A},\hat{B}]/2},$ (105)
we get:
$\begin{split}&\braket{n^{\prime}}{\hat{\mathcal{D}}^{\dagger}\left(\frac{q_{z}\ell_{B}}{\sqrt{2}}\right)}{n}=\\ &=e^{-|q_{z}|^{2}\ell_{B}^{2}/4}\braket{n^{\prime}}{\exp{\left(-\frac{q_{z}\ell_{B}\hat{a}^{\dagger}}{\sqrt{2}}\right)}\exp{\left(\frac{{q}^{*}_{z}\ell_{B}\hat{a}}{\sqrt{2}}\right)}}{n}~{}.\end{split}$ (106)
We further define:
${G}_{n^{\prime},n}({q}_{z}\ell_{B})\equiv\braket{n^{\prime}}{\exp{\left(-\frac{q_{z}\ell_{B}\hat{a}^{\dagger}}{\sqrt{2}}\right)}\exp{\left(\frac{{q}^{*}_{z}\ell_{B}\hat{a}}{\sqrt{2}}\right)}}{n}~{}.$ (107)
In the case were $n^{\prime}>n$, ${G}_{n^{\prime},n}({q}_{z}\ell_{B})$ can be calculated as:
$G_{n^{\prime}n}(q_{z}\ell_{B})=\sqrt{\frac{n!}{n^{\prime}!}}\bigg{(}\frac{\mathrm{i}q_{z}\ell}{\sqrt{2}}\bigg{)}^{\!n^{\prime}-n}L^{n^{\prime}-n}_{n}\bigg{(}\frac{\lvert q_{z}\rvert^{2}\ell_{B}^{2}}{2}\bigg{)}~{},$ (108)
where $L^{m}_{n}\left(X\right)$ is a Laguerre polynomial defined as:
$L^{m}_{n}\left(X\right)=\sum_{k=0}^{n}\binom{n+m}{n-k}\left(-X\right)^{\!k}~{}.$ (109)
The case $n^{\prime}<n$ can be obtained by noticing that
$G^{*}_{n^{\prime}n}(X)=G_{nn^{\prime}}(X)~{}.$ (110)
The matrix element of the density operator reads
$\begin{split}&\braket{k_{x}^{\prime},n^{\prime},\lambda^{\prime}}{\hat{n}_{\bm{q}_{\parallel}}}{k_{x},n,\lambda}=\\ &=e^{-\lvert q_{\parallel}\rvert^{2}\ell_{B}^{2}/4}\delta_{k_{x}^{\prime},k_{x}-q_{x}}\mathcal{F}_{n^{\prime},\lambda^{\prime},n,\lambda}(\bm{q}_{\parallel})~{},\end{split}$ (111)
where
$\begin{split}&\mathcal{F}_{n^{\prime},\lambda^{\prime},n,\lambda}(\bm{q}_{\parallel})\equiv\\ &\equiv C^{-}_{n^{\prime}}C^{-}_{n}G_{n^{\prime}-1,n-1}(\bar{q}_{z}\ell_{B})+\lambda\lambda^{\prime}C^{+}_{n^{\prime}}C^{+}_{n}G_{n^{\prime}n}(\bar{q}_{z}\ell_{B})~{}.\end{split}$ (112)
We can thus calculate the density–density response function as
$\chi^{(0)}_{nn}(q_{\parallel},\omega)=\sum_{n,\lambda,n^{\prime},\lambda^{\prime}}\frac{(f_{n\lambda}-f_{n^{\prime}\lambda^{\prime}})\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})}{\hbar\omega+\varepsilon_{\lambda,n}-\varepsilon_{\lambda^{\prime},n^{\prime}}+\mathrm{i}\eta}~{}.$ (113)
where $\varepsilon_{\lambda,n}\equiv\lambda(\hbar v_{\rm F}/\ell_{B})\sqrt{2n}$ and $\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})$ represents the squared modulus of the density matrix element,
$\begin{split}&\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})\equiv\sum_{k^{\prime}_{x},k_{x}}\lvert\braket{k_{x}^{\prime},n^{\prime},\lambda^{\prime}}{\hat{n}_{\bm{q}_{\parallel}}}{k_{x},n,\lambda}\rvert^{2}~{},\\ &=\frac{g}{2\pi\ell_{B}^{2}}e^{-q_{\parallel}^{2}\ell_{B}^{2}/2}|\mathcal{F}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})|^{2}~{},\end{split}$ (114)
where $|\mathcal{F}_{n^{\prime},\lambda^{\prime},n,\lambda}(q_{\parallel})|^{2}$ depends only on the modulus $q_{\parallel}$ of $\bm{q}_{\parallel}$ and reads
$|\mathcal{F}_{n^{\prime},\lambda^{\prime},n,\lambda}(q_{\parallel})|^{2}=\bigg{(}\frac{q_{\parallel}^{2}\ell_{B}^{2}}{2}\bigg{)}^{\!n_{>}-n_{<}}\Big{|}C^{-}_{n^{\prime}-1}C^{-}_{n-1}\sqrt{\frac{(n_{<}-1)!}{(n_{>}-1)!}}L^{n_{>}-n_{<}}_{n_{<}-1}\bigg{(}\frac{q_{\parallel}^{2}\ell_{B}^{2}}{2}\bigg{)}+\lambda\lambda^{\prime}C^{+}_{n^{\prime}}C^{+}_{n}\sqrt{\frac{n_{<}!}{n_{>}!}}L^{n_{>}-n_{<}}_{n_{<}}\bigg{(}\frac{q_{\parallel}^{2}\ell_{B}^{2}}{2}\bigg{)}\Big{|}^{2}~{},$ (115)
and $n_{>}={\rm max}(n^{\prime},n)$, $n_{<}={\rm min}(n^{\prime},n)$.
At $T=0$, the density–density response function is
$\chi^{(0)}_{nn}(q_{\parallel},\omega)=\sum_{n,\lambda\in\mathcal{N}_{\rm occ}}\sum_{n^{\prime},\lambda^{\prime}\in\mathcal{N}_{\rm unocc}}\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})\left(\frac{1}{\hbar\omega+\varepsilon_{\lambda,n}-\varepsilon_{\lambda^{\prime},n^{\prime}}+\mathrm{i}\eta}-\frac{1}{{\hbar\omega+\varepsilon_{\lambda^{\prime},n^{\prime}}-\varepsilon_{\lambda,n}+\mathrm{i}\eta}}\right)~{}.$ (116)
This concludes the calculation of the density–density response function $\chi^{(0)}_{nn}(q_{\parallel},\omega)$ for Landau levels in graphene.